### v1

# Quantum field theory treatment of neutrino spin-flavor oscillations in a magnetic field

Abstract
We study spin-flavor oscillations of neutrinos in a magnetic field within the quantum field theory approach in which neutrinos are virtual particles. Neutrinos are taken to be Majorana particles having a nonzero transition magnetic moment. We derive the dressed propagators of the neutrino mass eigenstates exactly accounting for the magnetic field contribution. The matrix element and the transition probability for spin-flavor oscillations are obtained. The leading term in the transition probability is shown to coincide with the result of the standard quantum mechanical description of neutrino oscillations. We also derive the correction to the transition probability which has the quantum field theory origin.
1 Introduction
The interaction of neutrinos with external fields can significantly modify the dynamics of neutrino oscillations. From the point of view of the phenomenology, the most important external field, contributing to flavor oscillations, is the electroweak interaction of neutrinos with background matter. It can result in the sizable amplification of the transition probability, known as the Mikheyev–Smirnov–Wolfenstein effect [1, 2], which is the most plausible explanation of the solar neutrino deficit [3].
Neutrinos are supposed to be electrically neutral particles. Nevertheless, these particles can interact with electromagnetic fields owing to nonzero magnetic moments which have purely anomalous origin (see, e.g., Ref. [4]). It should be noted that the electromagnetic properties of Dirac and Majorana neutrinos are completely different. Dirac neutrinos can have both diagonal and transition magnetic moments, whereas Majorana neutrinos can possess only transition ones. Despite the issue of the neutrino nature is still open, neutrinos are likely to be Majorana particles. At least, the smallness of the active neutrino masses is explained more naturally for Majorana neutrinos [5]. The recent advances in the neutrinos electromagnetic properties are reviewed in Ref. [6].
If neutrinos with transition magnetic moments interact with an external magnetic field, they undergo both flavor change and the spin flip. This process is called neutrino spin-flavor oscillations [7]. Neutrino spin-flavor oscillations were studied, e.g., in Ref. [8] to explain the possible anticorrelation between the solar activity and the solar neutrinos flux. Later, spin-flavor oscillations were understood in Ref. [9] to play a subdominant role for the solar neutrinos deficit problem. The description of spin-flavor oscillations of Majorana neutrinos expressed in terms of Weyl spinors was carried out in Ref. [10].
As a rule, neutrino oscillations in vacuum and in external fields are described within the quantum mechanical approach, which gives satisfactory results in the majority of situations. Nonetheless, this formalism has some shortcomings which were outlined in Ref. [11]. Numerous attempts to construct the description of neutrino flavor oscillations based on the quantum field theory (QFT) methods have been undertaken. Some of them are reviewed in Ref. [12].
We mention one of the QFT approaches which was put forward independently in Refs. [13, 14, 15]. Neutrinos are supposed to be virtual particles in this formalism. We also mention Refs. [16, 17], where QFT was applied to study neutrino flavor oscillations in external fields.
The approach in Refs. [13, 14, 15] allows one to treat neutrino oscillations in vacuum since it involves the propagators of the neutrino mass eigenstates. In vacuum, these propagators are diagonal in the neutrino types. If a neutrino interacts with an external field, the mass eigenstates propagators become nondiagonal. Finding such nondiagonal propagators in not a trivial task. A particular case of spin-flavor oscillations of Dirac neutrinos, possessing diagonal magnetic moments in the mass basis, was analyzed in frames of QFT in Ref. [18].
Recently, in Ref. [19], we developed a method which enables one to find the propagators of neutrino mass eigenstates in an external field. This approach is based on solving of the Dyson equations for propagators accounting for the interaction with external fields. Using this method in Ref. [19], we described neutrino flavor oscillations in background matter within QFT.
In the present work, we apply the results of Ref. [19] for the studies of neutrino spin-flavor oscillations in a magnetic field. Moreover, we consider the situation when the interaction with an external field is essentially nondiagonal in the mass basis. We shall see shortly in Sec. (2) that it is implemented for Majorana neutrinos. Therefore, the application of the results of Ref. [19] is crucial in this case. Note that the alternative formalism, based on relativistic quantum mechanics, for the description of neutrino flavor and spin-flavor oscillations in various external fields was considered in Ref. [20].
This work is organized in the following way. We remind briefly some basic issues of the Majorana neutrinos electrodynamics in Sec. 2. Then, in Sec. 3, we outline the QFT approach and adapt it for neutrino spin-flavor oscillations. We derive the dressed propagators for the neutrino mass eigenstates in a magnetic field by solving the corresponding Dyson equations in Sec. 4. The matrix element and the transition probability for neutrino spin-flavor oscillations are computed in Sec. 5. The results of the QFT and the standard quantum mechanical approaches are compared in Sec. 6. Finally, we conclude in Sec. 7. The propagators of massive Weyl neutrinos in vacuum are derived in Appendix.
2 Electrodynamics of Majorana neutrinos
In this section, we remind how massive Majorana neutrinos interact with electromagnetic fields.
Let us consider the system of two active flavor neutrinos $\nu_{\lambda}=(\nu_{e},\nu_{\mu})$. The mass matrix of these neutrinos is supposed to be nondiagonal. For simplicity, we take that no sterile neutrinos are present. To diagonalize the mass matrix we introduce the mass eigenstates $\psi_{a}$, $a=1,2$, by means of the matrix transformation,
$\nu_{\lambda}=\sum_{a}U_{\lambda a}\psi_{a},\quad(U_{\lambda a})=\left(\begin{array}[]{cc}\cos\theta&\sin\theta\\ -\sin\theta&\cos\theta\end{array}\right),$ (2.1)
where $\theta$ is the vacuum mixing angle. We assume that $\psi_{a}$ are Majorana particles having masses $m_{a}$.
We suppose that $\psi_{a}$ have the nonzero matrix of magnetic moments $(\mu_{ab})$. This matrix was found in Ref. [21] to be hermitian and purely imaginary. Hence, if we deal with two mass eigenstates, we can represent this matrix as $\mu_{ab}=\mu(\sigma_{2})_{ab}$, where $a,b=1,2$, $\sigma_{2}$ is the Pauli matrix, and $\mu$ is the transition magnetic moment. Then, we suppose that these neutrinos interact with an external electromagnetic field $F_{\mu\nu}=(\mathbf{E},\mathbf{B})$. The wave equations for the mass eigenstates can be formally obtained from the Dirac equation for $\psi_{a}$ [22],
$(\mathrm{i}\gamma^{\mu}\partial_{\mu}-m_{a})\psi_{a}-\frac{\mu_{ab}}{2}\sigma_{\mu\nu}F^{\mu\nu}\psi_{b}=0,$ (2.2)
where $\gamma^{\mu}$, $\gamma^{5}$, and $\sigma_{\mu\nu}=(\mathrm{i}/2)(\gamma_{\mu}\gamma_{\nu}-\gamma_{\nu}\gamma_{\mu})$ are the Dirac matrices. Since the mass eigenstates are taken to be Majorana particles, we represent them in terms of the Weyl spinors $\eta_{a}$, $\psi_{a}^{\mathrm{T}}=(\mathrm{i}\sigma_{2}\eta_{a}^{*},\eta_{a})$. Then, assuming that the electric field is absent and using Eq. (2.2), we derive the wave equations for the mass eigenstates,
$\mathrm{i}\dot{\eta}_{1}+(\bm{\sigma}\mathbf{p})\eta_{1}+\mathrm{i}m_{1}\sigma_{2}\eta_{1}^{*}-\mu(\bm{\sigma}\mathbf{B})\sigma_{2}\eta_{2}^{*}$ $=0,$
$\mathrm{i}\dot{\eta}_{2}+(\bm{\sigma}\mathbf{p})\eta_{2}+\mathrm{i}m_{2}\sigma_{2}\eta_{2}^{*}+\mu(\bm{\sigma}\mathbf{B})\sigma_{2}\eta_{1}^{*}$ $=0,$ (2.3)
where $\mathbf{p}=-\mathrm{i}\nabla$ is the momentum operator and $\bm{\sigma}=(\sigma_{1},\sigma_{2},\sigma_{3})$ are the Pauli matrices.
One can see in Eq. (2) that the neutrino magnetic interaction causes transitions between different mass eigenstates, $1\leftrightarrow 2$. Moreover, the magnetic interaction has the structure similar to the mass terms, which are known to induce neutrino-to-antineutrino oscillations [13, 23]. Indeed, both terms are $\propto\sigma_{2}\eta_{a,b}^{*}$. Thus, neutrinos undergo transitions $1\leftrightarrow\bar{2}$ and $\bar{1}\leftrightarrow 2$, where the bar denotes an antiparticle state, in an external magnetic field.
3 Quantum field theory formalism for neutrino oscillations
In this section, we briefly outline the quantum field theory based formalism for neutrino oscillations in which neutrinos are virtual particles [13, 14, 15]. We adapt this method for the description of spin-flavor oscillations.
Following Ref. [13], we assume that the source and the detector of neutrinos, separated by the distance $\mathbf{L}$, consist of heavy nuclei $N$. The production and detection of neutrinos is characterized by the interaction of these nuclei with charged leptons $l_{\alpha}^{-}$ and antileptons $l_{\beta}^{+}$, where $\alpha,\beta=e,\mu,\dotsc$. Thus, neutrino oscillations correspond to the process $l_{\beta}^{-}+N\to\text{neutrinos}\to l_{\alpha}^{+}+N^{\prime}$. If the flavor content of charged leptons changes, i.e. $\beta\neq\alpha$, one can speak about neutrino flavor oscillations. If electric charges of leptons also change, e.g., we have an electron in a source and antimuon in a detector, such a process is allowed in vacuum for Majorana neutrinos. It was connected to neutrino-to-antineutrino oscillations in Refs. [13, 23]. However, if we study the transition $l_{\beta}^{-}+N\to\text{neutrinos}\to l_{\alpha}^{+}+N^{\prime}$ in an external field which neutrinos interact with, it is more appropriate to call it neutrino spin-flavor oscillations. The considered process is schematically depicted in Fig. 1.
Figure 1: The schematic Feynman diagram for the reaction $l_{\beta}^{-}+N\to\text{neutrinos}\to l_{\alpha}^{+}+N^{\prime}$. The neutrino line has two counter-directional arrows since we deal with the Majorana particles. Moreover, the neutrino line is broad since the mass eigenstates interact with a magnetic field [see Eq. (2)]. Note that we take into account the magnetic interaction exactly (see Sec. 4).
The $S$-matrix element of the reaction in question reads [13]
$S=-\frac{1}{2}\left(\sqrt{2}G_{\mathrm{int}}\right)^{2}\int\mathrm{d}^{4}x\mathrm{d}^{4}y\Big{\langle}l_{\alpha}^{+},N,N^{\prime}\Big{|}T\left\{j^{\mu}(x)J_{\mu}(x)j^{\nu}(y)J_{\nu}(y)\right\}\Big{|}l_{\beta}^{-},N,N^{\prime}\Big{\rangle},$ (3.1)
where $G_{\mathrm{int}}$ is the coupling constant,
$j_{\mu}=\sum_{\lambda}\bar{\nu}_{\lambda}\gamma_{\mu}^{\mathrm{L}}l_{\lambda},$ (3.2)
is the leptons current, $\gamma_{\mu}^{\mathrm{L}}=\gamma_{\mu}(1-\gamma^{5})/2$, and $J_{\mu}(x)$ is the nuclear current operator. Then, we represent the operator valued wave function of leptons as
$l_{\lambda}(x)=\int\mathrm{d}^{3}p\left[a_{\lambda}(p)u_{\lambda}(p)e^{-\mathrm{i}p_{\lambda}x}+b_{\lambda}^{\dagger}(p)v_{\lambda}(p)e^{\mathrm{i}p_{\lambda}x}\right],$ (3.3)
where $b_{\lambda}^{\dagger}(p)$ and $a_{\lambda}(p)$ are the creation and annihilation operators of antileptons and leptons, $u_{\lambda}(p)$ and $v_{\lambda}(p)$ are the basis spinors, $p_{\lambda}^{\mu}=(E_{\lambda},\mathbf{p})$, $E_{\lambda}=\sqrt{\mathfrak{m}_{\lambda}^{2}+p^{2}}$ is the lepton energy, and $\mathfrak{m}_{\lambda}$ is its mass. We take that $\left|l_{\beta}^{-}\right\rangle=a_{\beta}^{\dagger}(p_{\beta})\left|0\right\rangle$ and $\left\langle l_{\alpha}^{+}\right|=\left\langle 0\right|b_{\alpha}(p_{\alpha})$, i.e. we deal with the plane waves incoming and outgoing leptons. Eventually, we rewrite Eq. (3.1) in the following way:
$S=$ $2G_{\mathrm{int}}^{2}\sum_{ab}U_{\alpha a}^{*}U_{\beta b}^{*}v_{\alpha}^{\mathrm{T}}(p_{\alpha})(\gamma_{\mathrm{L}}^{\mu})^{\mathrm{T}}$
×$\times\left(\int\mathrm{d}^{4}x\mathrm{d}^{4}ye^{\mathrm{i}p_{\alpha}x-\mathrm{i}p_{\beta}y}J_{\mu}(x)J_{\nu}(y)\big{\langle}0\big{|}T\left\{(\bar{\psi}_{a}(x))^{\mathrm{T}}\bar{\psi}_{b}(y)\right\}\big{|}0\big{\rangle}\right)\gamma_{\mathrm{L}}^{\nu}u_{\beta}(p_{\beta}),$ (3.4)
where we account for Eqs. (3.2) and (3.3).
We assume that the nuclei in the source and the detector are quite heavy. Hence, their positions, $\mathbf{x}_{1}$ and $\mathbf{x}_{2}$, are unchanged, with $\mathbf{x}_{2}-\mathbf{x}_{1}=\mathbf{L}$. In this case, the nuclear currents in Eq. (3) have the form,
$J_{\mu}(x_{0},\mathbf{x})\propto\delta_{\mu 0}\delta(\mathbf{x}-\mathbf{x}_{2}),\quad J_{\nu}(y_{0},\mathbf{y})\propto\delta_{\nu 0}\delta(\mathbf{y}-\mathbf{x}_{1}).$ (3.5)
We take that the neutrino mass eigenstates are Majorana particles, i.e. $\left(\psi_{a}\right)^{c}=C\bar{\psi}_{a}^{\mathrm{T}}=\psi_{a}$, where $C$ is the charge conjugation matrix. In this case,
$v_{\alpha}^{\mathrm{T}}(p_{\alpha})(\gamma_{\mathrm{L}}^{0})^{\mathrm{T}}\left\langle 0\left|T\left\{(\bar{\psi}_{a}(x))^{\mathrm{T}}\bar{\psi}_{b}(y)\right\}\right|0\right\rangle\gamma_{\mathrm{L}}^{0}u_{\beta}(p_{\beta})\\ =\kappa_{\alpha\mathrm{R}}^{\dagger}(p_{\alpha})\left\langle 0\left|T\left\{\xi_{a}(x)\eta_{b}^{\dagger}(y)\right\}\right|0\right\rangle\kappa_{\beta\mathrm{L}}(p_{\beta}).$ (3.6)
Here we use the representations
$\psi_{a}=\left(\begin{array}[]{c}\xi_{a}\\ \eta_{a}\end{array}\right),\quad u_{\alpha,\beta}(p_{\alpha,\beta})=\left(\begin{array}[]{c}\kappa_{\alpha,\beta\mathrm{R}}(p_{\alpha,\beta})\\ \kappa_{\alpha,\beta\mathrm{L}}(p_{\alpha,\beta})\end{array}\right).$ (3.7)
Note that we keep $\xi_{a}$ instead of $\mathrm{i}\sigma_{2}\eta_{a}^{*}$ in Eq. (3.7). We also use the known expressions, $\gamma_{\mathrm{R}}^{\mu}=C(\gamma_{\mathrm{L}}^{\mu})^{\mathrm{T}}C$ and $v_{\alpha}(p_{\alpha})=\left(u_{\alpha}(p_{\alpha})\right)^{c}$, to derive Eq. (3.6).
Using Eq. (3.6) and performing the integrations as in Ref. [13], we represent Eq. (3) in the form,
$S=4\pi G_{\mathrm{int}}^{2}e^{-i\mathbf{p}_{\alpha}\mathbf{x}_{2}+i\mathbf{p}_{\beta}\mathbf{x}_{1}}\delta(E_{\alpha}-E_{\beta})\mathcal{M}_{\beta\to\bar{\alpha}},$ (3.8)
where the matrix element is
$\mathcal{M}_{\beta\to\bar{\alpha}}=\sum_{ab}U_{\alpha a}^{*}U_{\beta b}^{*}\kappa_{\alpha+}^{\dagger}(p_{\alpha})\left(\int\frac{\mathrm{d}^{3}q}{(2\pi)^{3}}\Sigma_{ab}(E,\mathbf{q})e^{\mathrm{i}\mathbf{qL}}\right)\kappa_{\beta-}(p_{\beta}),$ (3.9)
where $E=(E_{\alpha}+E_{\beta})/2$ and $\Sigma_{ab}(q_{0},\mathbf{q})$ are the 4D Fourier images of the nondiagonal propagators,
$\Sigma_{ab}(x-y)=\left\langle 0\left|T\left\{\xi_{a}(x)\eta_{b}^{\dagger}(y)\right\}\right|0\right\rangle,$ (3.10)
We consider utrarelativistic incoming and outgoing leptons. Therefore, in Eq. (3.9), we replace $\kappa_{\beta\mathrm{L}}$ and $\kappa_{\alpha\mathrm{R}}$ with the helicity states $\kappa_{\beta-}$ and $\kappa_{\alpha+}$.
The probability of neutrino spin-flavor oscillations is $P_{\beta\to\bar{\alpha}}\propto|\mathcal{M}_{\beta\to\bar{\alpha}}|^{2}$. In case of the Majorana neutrinos motion in vacuum, the propagators are diagonal, $\Sigma_{ab}\propto\delta_{ab}$. In this situation, the process $\beta\to\bar{\alpha}$ corresponds to neutrino-to-antineutrino oscillations, as we mentioned earlier. The probability of $\nu_{\alpha}\to\bar{\nu}_{\beta}$ oscillations was found in Refs. [13, 23] to be of the order $\sim(m_{1,2}/E)^{2}$.
If Majorana neutrinos have a transition magnetic moment, as one can see in Eq. (2), the wave equations for $\eta_{1}$ and $\eta_{2}$ are coupled by the magnetic interaction. This interaction induces the transitions between $\eta_{1,2}$ and $\mathrm{i}\sigma_{2}\eta_{2,1}^{*}=\xi_{2,1}$. It means that both diagonal, $\Sigma_{11}$ and $\Sigma_{22}$, and nondiagonal, $\Sigma_{12}$ and $\Sigma_{21}$, propagators in Eq. (3.10) can contribute to the transition probability. In case we study spin-flavor oscillations like $\nu_{e}\to\bar{\nu}_{\mu}$, the matrix element in Eq. (3.9) takes the form,
$\mathcal{M}_{e\to\bar{\mu}}=\kappa_{\mu+}^{\dagger}(p_{\mu})\int\frac{\mathrm{d}^{3}q}{(2\pi)^{3}}e^{\mathrm{i}\mathbf{qL}}\left[\frac{1}{2}\left(\Sigma_{22}-\Sigma_{11}\right)\sin 2\theta+\cos^{2}\theta\Sigma_{21}-\sin^{2}\theta\Sigma_{12}\right]\kappa_{e-}(p_{e}),$ (3.11)
where we use Eq. (2.1). Our main goal is to find the propagators $\Sigma_{ab}$ in Eq. (3.11) in the presence of the magnetic field.
4 Propagators of massive Majorana neutrinos in a magnetic field
In this section, we derive both diagonal and nondiagonal propagators of Majorana neutrinos interacting with a magnetic field. For this purpose, we use the method developed in Ref. [19], where the propagators in question are the solutions of the Dyson equations exactly accounting for the interaction with an external field. First, we derive the Dyson equations within the perturbative approach, with all the terms in the perturbation series being summed up. Then, the Dyson equations are solved exactly for ultrarelativistic neutrinos. We deal with two mass eigenstates since we mentioned in Ref. [19] that extension of the formalism for the greater number of neutrinos is ambiguous.
The magnetic interaction in Eq. (2) mixes both opposite helicities and different mass eigenstates. Thus, if we have a left-handed lepton in a source and a right-handed one in a detector, like in Eq. (3.11), the dressed propagators are the sums of the following perturbative series:
$-\mathrm{i}\Sigma_{12}=$ $(-\mathrm{i}S_{1\mathrm{R}})V(-\mathrm{i}S_{2\mathrm{L}})+(-\mathrm{i}S_{1\mathrm{R}})V(-\mathrm{i}S_{2\mathrm{L}})(-V)(-\mathrm{i}S_{1\mathrm{R}})V(-\mathrm{i}S_{2\mathrm{L}})+\dotsb,$
$-\mathrm{i}\Sigma_{21}=$ $(-\mathrm{i}S_{2\mathrm{R}})(-V)(-\mathrm{i}S_{1\mathrm{L}})+(-\mathrm{i}S_{2\mathrm{R}})(-V)(-\mathrm{i}S_{1\mathrm{L}})V(-\mathrm{i}S_{2\mathrm{R}})(-V)(-\mathrm{i}S_{1\mathrm{L}})+\dotsb,$
$-\mathrm{i}\Sigma_{11}=$ $-\mathrm{i}S_{1\mathrm{L}}+(-\mathrm{i}S_{1\mathrm{L}})V(-\mathrm{i}S_{2\mathrm{R}})(-V)(-\mathrm{i}S_{1\mathrm{L}})+\dotsb,$
$-\mathrm{i}\Sigma_{22}=$ $-\mathrm{i}S_{2\mathrm{L}}+(-\mathrm{i}S_{2\mathrm{L}})(-V)(-\mathrm{i}S_{1\mathrm{R}})V(-\mathrm{i}S_{2\mathrm{L}})+\dotsb,$ (4.1)
where $V=\mathrm{i}\mu(\bm{\sigma}\mathbf{B})$ is the magnetic interaction and $S_{a\mathrm{L,R}}$ are the propagators of left-handed neutrinos and right-handed antineutrinos. Since the magnetic interaction contributes to the transitions $a\leftrightarrow\bar{b}$ only, the diagonal propagators $S_{a\mathrm{L,R}}$ should be taken in vacuum. We give them in Eqs. (A.2) and (A.10). We also take into account in Eq. (4) that the magnetic interaction contributes to the wave equations for $\eta_{1}$ and $\eta_{2}$ with opposite signs; cf. Eq. (2). The perturbative series in Eq. (4) are illustrated in Fig. 2.
Figure 2: Feynman diagrams corresponding to Eq. (4). 4D Fourier images of the bare propagators $-\mathrm{i}S_{a\mathrm{R,L}}$, given in Eqs. (A.2) and (A.10), are depicted by thin solid lines [24]. The broad solid lines correspond to the dressed propagators $-\mathrm{i}\Sigma_{ab}$. The potential of the magnetic interaction is $V=\mathrm{i}\mu(\bm{\sigma}\mathbf{B})$.
The series in Eq. (4) can be summed up. The result is
$\mathrm{i}\Sigma_{12}^{-1}$ $=-(S_{1\mathrm{R}}VS_{2\mathrm{L}})^{-1}+V,\quad\mathrm{i}\Sigma_{21}^{-1}=(S_{2\mathrm{R}}VS_{1\mathrm{L}})^{-1}-V,$
$\Sigma_{11}^{-1}$ $=S_{1\mathrm{L}}^{-1}-VS_{2\mathrm{R}}V,\quad\Sigma_{22}^{-1}=S_{2\mathrm{L}}^{-1}-VS_{1\mathrm{R}}V.$ (4.2)
We notice the diagonal propagators in Eq. (4) do not contribute to the matrix element since $\mathcal{M}_{e\to\bar{\mu}}$ is averaged in Eq. (3.11) over the leptonic states with opposite helicities. Hence, we should look for $\Sigma_{12}$ and $\Sigma_{21}$ only.
One can see in the expressions for $S_{a\mathrm{L,R}}$ in Eqs. (A.2) and (A.10) that these propagators contain the projectors $(1\mp\bm{\sigma}\hat{p})$ for ultrarelativistic neutrinos. Thus, the inverse operators $S_{a\mathrm{L,R}}^{-1}$, which are present in Eq. (4), do not exist for such neutrinos. However, we keep the ratio $\varsigma_{a}=p/E_{a}$ in soling of Eq. (4) and set $\varsigma_{a}\to 1$ in the final result. To proceed we choose the geometry of the system. Let us take that $\mathbf{L}=L\mathbf{e}_{z}$, where $\mathbf{e}_{z}$ is the unit vector along the $z$-axis. We also assume that the process is considered in the forward scattering approximation, i.e. $\mathbf{p}_{e}\upuparrows\mathbf{p}_{\bar{\mu}}\upuparrows\mathbf{e}_{z}$, which is valid for ultrarelativistic leptons. The magnetic field is taken to be transverse to the charged leptons motion, $\mathbf{B}=B\mathbf{e}_{x}$, where $\mathbf{e}_{x}$ is the unit vector along the $x$-axis.
We provide the detailed calculations for $\Sigma_{12}$ only. Using Eqs. (A.2) and (A.10), the exact solution of the Dyson equation for $\Sigma_{12}$ in Eq. (4) has the form,
$\Sigma_{12}(E,\mathbf{q})=$ $-\frac{1}{4\alpha\varsigma_{1}\varsigma_{2}\mu^{2}(\hat{q}\mathbf{B})^{2}+(\mu B)^{2}[\alpha^{2}(1-\varsigma_{1}^{2})(1-\varsigma_{2}^{2})-2\alpha(1+\varsigma_{1}\varsigma_{2})+1]}$
×$\times\big{\{}\alpha(\varsigma_{1}-\varsigma_{2})\mu(\hat{q}\mathbf{B})+(\bm{\sigma}\cdot[\mathrm{i}\alpha(\varsigma_{1}+\varsigma_{2})\mu(\hat{q}\times\mathbf{B})-2\alpha\varsigma_{1}\varsigma_{2}\mu(\hat{q}\mathbf{B})\hat{q}$
$+\mu\mathbf{B}(\alpha(1+\varsigma_{1}\varsigma_{2})-1)])\big{\}},$ (4.3)
where
$\alpha=\frac{4(E-E_{1}+\mathrm{i}0)(E-E_{2}+\mathrm{i}0)}{\mu^{2}B^{2}(1-\varsigma_{1}^{2})(1-\varsigma_{2}^{2})},$ (4.4)
and $E_{1,2}=\sqrt{q^{2}+m_{1,2}^{2}}$. For ultrarelativistic neutrinos, one has that $\varsigma_{1,2}\to 1$ and Eq. (4) takes the form,
$\Sigma_{12}=-\frac{\mu(\bm{\sigma}\cdot[\mathbf{B}+\mathrm{i}(\hat{q}\times\mathbf{B})-(\hat{q}\mathbf{B})\hat{q}])}{2[(E-E_{1}+\mathrm{i}0)(E-E_{2}+\mathrm{i}0)-\mu^{2}B^{2}+\mu^{2}(\hat{q}\mathbf{B})^{2}]}.$ (4.5)
Analogously to Eq. (4.5), the rest of the propagators, resulting from Eq. (4), can be derived. We present them for ultrarelativistic neutrinos,
$\Sigma_{11}$ $=\frac{\mathrm{i}(E-E_{2})[1-(\bm{\sigma}\hat{q})]}{2[(E-E_{1}+\mathrm{i}0)(E-E_{2}+\mathrm{i}0)-\mu^{2}B^{2}+\mu^{2}(\hat{q}\mathbf{B})^{2}]},$
$\Sigma_{22}$ $=\frac{\mathrm{i}(E-E_{1})[1-(\bm{\sigma}\hat{q})]}{2[(E-E_{1}+\mathrm{i}0)(E-E_{2}+\mathrm{i}0)-\mu^{2}B^{2}+\mu^{2}(\hat{q}\mathbf{B})^{2}]},$ (4.6)
and $\Sigma_{21}=-\Sigma_{12}$.
5 Transition probability of spin-flavor oscillations
In this section, we calculate the 3D Fourier transform of the dressed propagators derived in Sec. 4, which the matrix element depends on; cf. Eq. (3.11). Then, we compute the transition probability of $\nu_{e}\to\bar{\nu}_{\mu}$ oscillations.
We present the detailed calculation of the $\Sigma_{12}$ contribution to the matrix element. Following Ref. [19], we use the cylindrical coordinates $(\rho,\phi,z)$ to compute the integral in Eq. (3.11), $\mathbf{q}=\rho(\cos\phi\mathbf{e}_{x}+\sin\phi\mathbf{e}_{y})+z\mathbf{e}_{z}$. The spinors of the charged leptons are $\kappa_{e-}^{\mathrm{T}}=(0,1)$ and $\kappa_{\mu+}^{\mathrm{T}}=(1,0)$. Dropping the terms which vanish in the integration over $\phi$, as well as averaging over the leptons states, one gets that
$\kappa_{\mu+}^{\dagger}\Sigma_{12}\kappa_{e-}=$ $\frac{f}{[\mathcal{E}_{+}(\mathbf{q})-\mathcal{E}_{-}(\mathbf{q})]}\left(1+\frac{z}{\sqrt{\rho^{2}+z^{2}}}-\frac{\rho^{2}\cos^{2}\phi}{\rho^{2}+z^{2}}\right)$
×$\times\left(\frac{1}{E-\mathcal{E}_{+}(\mathbf{q})+\mathrm{i}0}-\frac{1}{E-\mathcal{E}_{-}(\mathbf{q})+\mathrm{i}0}\right),$ (5.1)
where
$f(\rho,\phi,z)=-\frac{\mu B}{2}\left(1+\frac{z}{\sqrt{\rho^{2}+z^{2}}}-\frac{\rho^{2}\cos^{2}\phi}{\rho^{2}+z^{2}}\right),$ (5.2)
and
±$\mathcal{E}_{\pm}(\mathbf{q})=\bar{E}\pm\sqrt{\Delta E^{2}+(\mu B)^{2}\left(1-\frac{\rho^{2}\cos^{2}\phi}{\rho^{2}+z^{2}}\right)}.$ (5.3)
Here $\bar{E}=[E_{2}(\mathbf{q})+E_{1}(\mathbf{q})]/2$ and $\Delta E=[E_{2}(\mathbf{q})-E_{1}(\mathbf{q})]/2$.
Let us compute one of the integrals, which contributes to Eq. (5). Namely
$I_{+}=$ $\frac{1}{(2\pi)^{3}}\int_{0}^{2\pi}\mathrm{d}\phi\int_{0}^{\infty}\rho\mathrm{d}\rho\int_{-\infty}^{+\infty}\mathrm{d}z\frac{fe^{\mathrm{i}zL}}{(\mathcal{E}_{+}-\mathcal{E}_{-})(E-\mathcal{E}_{+}+\mathrm{i}0)}$
$=-\frac{\mathrm{i}}{(2\pi)^{2}}\int_{0}^{2\pi}\mathrm{d}\phi\sum_{\left\{z_{+}\right\}}\int_{0}^{\infty}\rho\mathrm{d}\rho\left.\frac{f(\rho,\phi,z_{+})e^{\mathrm{i}z_{+}L}}{(\mathcal{E}_{+}-\mathcal{E}_{-})\frac{\mathrm{d}\mathcal{E}_{+}}{\mathrm{d}z}}\right|_{z=z_{+}},$ (5.4)
[…]
We define the new quantity
$\mathcal{E}_{+}(E)=E+\sqrt{\left(\frac{\Delta m^{2}}{4E}\right)^{2}+(\mu B)^{2}\left(1-\frac{\rho^{2}}{E^{2}}\cos^{2}\phi\right)},$ (5.5)
where $\Delta m^{2}=m_{2}^{2}-m_{1}^{2}>0$. Note that $\mathcal{E}_{+}(E)\neq\mathcal{E}_{+}(\mathbf{q})$; cf. Eq. (5.3). We can approximately set that $\left.\frac{\mathrm{d}\mathcal{E}_{+}(\mathbf{q})}{\mathrm{d}z}\right|{}_{z=z_{+}}\approx\frac{z_{+}}{E}$ and $\mathcal{E}_{+}^{2}(E)\approx\rho^{2}+z_{+}^{2}$. Moreover, for ultrarelativistic leptons, as well as neutrinos, having $E\gg\mathfrak{E}$ and a rather weak magnetic field, obeying $(\mu B)^{2}\ll E\mathfrak{E}$, one has that
$\mathcal{E}_{+}^{2}(E)\approx[\mathcal{E}_{+}^{(0)}(E)]^{2}-\frac{(\mu B)^{2}}{E\mathfrak{E}}\rho^{2}\cos^{2}\phi,$ (5.6)
where
$\mathcal{E}_{+}^{(0)}(E)=E+\mathfrak{E},\quad\mathfrak{E}=\sqrt{(\mu B)^{2}+\left(\frac{\Delta m^{2}}{4E}\right)^{2}}.$ (5.7)
Using Eq. (5.6), one gets that the poles $z_{+}$ obey the relation,
$[\mathcal{E}_{+}^{(0)}(E)]^{2}=\rho^{2}\left(1+\frac{(\mu B)^{2}}{E\mathfrak{E}}\cos^{2}\phi\right)+z_{+}^{2}.$ (5.8)
Finally, we obtain the following expressions for the poles in the explicit form:
$z_{+}$ $=\sqrt{1+\frac{(\mu B)^{2}}{E\mathfrak{E}}\cos^{2}\phi}\times\begin{cases}\sqrt{\rho_{0}^{2}-\rho^{2}},&\text{if}\ \rho<\rho_{0},\\ \mathrm{i}\sqrt{\rho^{2}-\rho_{0}^{2}},&\text{if}\ \rho>\rho_{0},\end{cases}$ (5.9)
where $\rho_{0}=(E+\mathfrak{E})/\sqrt{1+\frac{(\mu B)^{2}}{E\mathfrak{E}}\cos^{2}\phi}$.
Using Eqs. (5.2) and (5.9), we rewrite Eq. (5) as
$I_{+}=$ $\frac{\mathrm{i}\mu BE}{16\pi^{2}\mathfrak{E}}\int_{0}^{2\pi}\frac{\mathrm{d}\phi}{\sqrt{1+\frac{(\mu B)^{2}}{E\mathfrak{E}}\cos^{2}\phi}}\Bigg{[}\int_{0}^{\rho_{0}}\frac{\rho\mathrm{d}\rho\exp\left(\mathrm{i}\tilde{L}\sqrt{\rho_{0}^{2}-\rho^{2}}\right)}{\sqrt{\rho_{0}^{2}-\rho^{2}}}$
×$\times\left(1+\frac{\sqrt{\rho_{0}^{2}-\rho^{2}}}{\tilde{E}}-\frac{\rho^{2}\cos^{2}\phi}{E^{2}}\left[1-\frac{(\mu B)^{2}}{2\mathfrak{E}^{2}}\right]\right)$
$-\mathrm{i}\int_{\rho_{0}}^{\infty}\frac{\rho\mathrm{d}\rho\exp\left(-\tilde{L}\sqrt{\rho^{2}-\rho_{0}^{2}}\right)}{\sqrt{\rho^{2}-\rho_{0}^{2}}}\left(1+\mathrm{i}\frac{\sqrt{\rho^{2}-\rho_{0}^{2}}}{\tilde{E}}-\frac{\rho^{2}\cos^{2}\phi}{E^{2}}\left[1-\frac{(\mu B)^{2}}{2\mathfrak{E}^{2}}\right]\right)\Bigg{]},$ (5.10)
where
$\tilde{L}=L\sqrt{1+\frac{(\mu B)^{2}}{E\mathfrak{E}}\cos^{2}\phi},\quad\tilde{E}=\frac{E}{\sqrt{1+\frac{(\mu B)^{2}}{E\mathfrak{E}}\cos^{2}\phi}}.$ (5.11)
Making the integrations over $\rho$, we represent Eq. (5) in the form,
$I_{+}\approx$ $\frac{\mu BEe^{i(E+\mathfrak{E})}}{8\pi{}^{2}\mathfrak{E}L}\left(1+\frac{\mathfrak{E}}{2E}\right)\int_{0}^{2\pi}\frac{\mathrm{d}\phi}{1+\frac{(\mu B)^{2}}{E\mathfrak{E}}\cos^{2}\phi}$
$\approx\frac{\mu BEe^{\mathrm{i}(E+\mathfrak{E})L}}{4\pi\mathfrak{E}L}\left(1+\frac{\left(\frac{\Delta m^{2}}{4E}\right)^{2}}{2\mathfrak{E}E}\right),$ (5.12)
where we assume that the distance between the source and the detector is great, $L\gg E^{-1}$.
Analogously to Eq. (5), one computes the integral,
$I_{-}=\int\frac{\mathrm{d}^{3}q}{(2\pi)^{3}}\frac{fe^{\mathrm{i}\mathbf{qL}}}{(\mathcal{E}_{+}-\mathcal{E}_{-})(E-\mathcal{E}_{-}+\mathrm{i}0)}.$ (5.13)
We give just the final result for $I_{-}$ in Eq. (5.13),
$I_{-}\approx\frac{\mu BEe^{\mathrm{i}(E-\mathfrak{E})L}}{4\pi\mathfrak{E}L}\left(1-\frac{\left(\frac{\Delta m^{2}}{4E}\right)^{2}}{2\mathfrak{E}E}\right).$ (5.14)
The contributions of the diagonal propagators $\Sigma_{aa}$ in Eq. (4) to the matrix element $\mathcal{M}_{e\to\bar{\mu}}$ are vanishing.
Since $\Sigma_{21}=-\Sigma_{12}$, one gets from Eqs. (3.11), (5.11), and (5.14) that
$\mathcal{M}_{e\to\bar{\mu}}=I_{-}-I_{+}=-\frac{\mathrm{i}\mu BEe^{\mathrm{i}EL}}{2\pi\mathfrak{E}L}\left(\sin\mathfrak{E}L-\frac{\mathrm{i}\left(\Delta m^{2}\right)^{2}}{32\mathfrak{E}E^{3}}\cos\mathfrak{E}L\right).$ (5.15)
The probability of the process $\nu_{e}\to\bar{\nu}_{\mu}$ is $P_{\nu_{e}\to\bar{\nu}_{\mu}}\propto|\mathcal{M}_{e\to\bar{\mu}}|^{2}$. Based on Eq. (5.15) and accounting for Eq. (5.7), we obtain that $P_{\nu_{e}\to\bar{\nu}_{\mu}}=P_{\nu_{e}\to\bar{\nu}_{\mu}}^{(0)}+\delta P_{\nu_{e}\to\bar{\nu}_{\mu}}$, where
$P_{\nu_{e}\to\bar{\nu}_{\mu}}^{(0)}(L)$ $=\frac{(\mu B)^{2}}{(\mu B)^{2}+\left(\frac{\Delta m^{2}}{4E}\right)^{2}}\sin^{2}\left(\sqrt{(\mu B)^{2}+\left(\frac{\Delta m^{2}}{4E}\right)^{2}}L\right),$ (5.16)
$\delta P_{\nu_{e}\to\bar{\nu}_{\mu}}(L)$ $=\frac{(\mu B)^{2}(\Delta m^{2})^{4}}{4096E^{6}\left[(\mu B)^{2}+\left(\frac{\Delta m^{2}}{4E}\right)^{2}\right]^{2}}\sin^{2}\left(2\sqrt{(\mu B)^{2}+\left(\frac{\Delta m^{2}}{4E}\right)^{2}}L\right).$ (5.17)
We compare the transition probability in Eq. (5.17) with the result of the quantum mechanical approach shortly in Sec. 6.
6 Quantum mechanical description of spin-flavor oscillations
In this section, we analyze the quantum mechanical description of spin-flavor oscillations of Majorana neutrinos in an external magnetic field. This study can be done in both flavor and mass bases. We take the flavor basis since it is more common in the standard quantum mechanical approach.
We suppose that two flavor neutrinos, $\nu_{e}$ and $\nu_{\mu}$, propagate along the $z$-axis and interact with the transverse external magnetic field $\mathbf{B}=(B,0,0)$, as in Sec. 4. Despite one can reveal the nature of neutrinos only in the mass basis, we can still assume that $\nu_{\lambda}$, $\lambda=e,\mu$, are Majorana particles. The interaction of these neutrinos with the magnetic field is owing to the transition magnetic moment $\mu$. The effective Schrodinger equation in the basis $\nu^{\mathrm{T}}=(\nu_{e},\nu_{\mu},\bar{\nu}_{e},\bar{\nu}_{\mu})$ was obtained in Ref. [7],
$\mathrm{i}\dot{\nu}=H_{f}\nu,\quad H_{f}=\left(\begin{array}[]{cccc}-\frac{\Delta m^{2}}{4E_{\nu}}\cos 2\theta&\frac{\Delta m^{2}}{4E_{\nu}}\sin 2\theta&0&\mathrm{i}\mu B\\ \frac{\Delta m^{2}}{4E_{\nu}}\sin 2\theta&\frac{\Delta m^{2}}{4E_{\nu}}\cos 2\theta&-\mathrm{i}\mu B&0\\ 0&\mathrm{i}\mu B&-\frac{\Delta m^{2}}{4E_{\nu}}\cos 2\theta&\frac{\Delta m^{2}}{4E_{\nu}}\sin 2\theta\\ -\mathrm{i}\mu B&0&\frac{\Delta m^{2}}{4E_{\nu}}\sin 2\theta&\frac{\Delta m^{2}}{4E_{\nu}}\cos 2\theta\end{array}\right),$ (6.1)
where $E_{\nu}$ is the mean energy of neutrinos, and $\theta$ is the vacuum mixing angle defined in Eq. 2.1. We already mentioned in Sec. 2 (see also Ref. [21]) that the magnetic moments matrix $(\mu_{ab})=\mu\sigma_{2}$ is hermitian and purely imaginary in the mass basis. Taking into account Eq. 2.1, we get that this matrix in the flavor basis reads $(M_{\lambda\lambda^{\prime}})=U(\mu_{ab})U^{\dagger}=\mu\sigma_{2}$, that is used in Eq. (6.1).
The general solution of Eq. (6.1), which satisfies the given initial condition $\nu_{0}$, has the form,
$\nu(t)=\left(\left[\left(U_{1}\otimes U_{1}^{\dagger}\right)+\left(U_{2}\otimes U_{2}^{\dagger}\right)\right]e^{-\mathrm{i}\mathfrak{\mathfrak{E}_{\nu}}t}+\left[\left(V_{1}\otimes V_{1}^{\dagger}\right)+\left(V_{2}\otimes V_{2}^{\dagger}\right)\right]e^{\mathrm{i}\mathfrak{E}_{\nu}t}\right)\nu_{0},$ (6.2)
where
$\mathfrak{E}_{\nu}=\sqrt{\left(\frac{\Delta m^{2}}{4E_{\nu}}\right)^{2}+(\mu B)^{2}},$ (6.3)
and
$U_{1}$ $=\frac{1}{\sqrt{2\mathfrak{E}_{\nu}}}\left(\begin{array}[]{c}\frac{\mathrm{i}\mu B}{\sqrt{\mathfrak{E}_{\nu}+\frac{\Delta m^{2}}{4E_{\nu}}\cos 2\theta}}\\ 0\\ \frac{\frac{\Delta m^{2}}{4E_{\nu}}\sin 2\theta}{\sqrt{\mathfrak{E}_{\nu}+\frac{\Delta m^{2}}{4E_{\nu}}\cos 2\theta}}\\ \sqrt{\mathfrak{E}_{\nu}+\frac{\Delta m^{2}}{4E_{\nu}}\cos 2\theta}\end{array}\right),\quad V_{1}=\frac{1}{\sqrt{2\mathfrak{E}_{\nu}}}\left(\begin{array}[]{c}\sqrt{\mathfrak{E}_{\nu}+\frac{\Delta m^{2}}{4E_{\nu}}\cos 2\theta}\\ -\frac{\frac{\Delta m^{2}}{4E_{\nu}}\sin 2\theta}{\sqrt{\mathfrak{E}_{\nu}+\frac{\Delta m^{2}}{4E_{\nu}}\cos 2\theta}}\\ 0\\ \frac{\mathrm{i}\mu B}{\sqrt{\mathfrak{E}_{\nu}+\frac{\Delta m^{2}}{4E_{\nu}}\cos 2\theta}}\end{array}\right),$ (6.12)
$U_{2}$ $=\frac{1}{\sqrt{2\mathfrak{E}_{\nu}}}\left(\begin{array}[]{c}\frac{\frac{\Delta m^{2}}{4E_{\nu}}\sin 2\theta}{\sqrt{\mathfrak{E}_{\nu}+\frac{\Delta m^{2}}{4E_{\nu}}\cos 2\theta}}\\ \sqrt{\mathfrak{E}_{\nu}+\frac{\Delta m^{2}}{4E_{\nu}}\cos 2\theta}\\ \frac{\mathrm{i}\mu B}{\sqrt{\mathfrak{E}_{\nu}+\frac{\Delta m^{2}}{4E_{\nu}}\cos 2\theta}}\\ 0\end{array}\right),\quad V_{2}=\frac{1}{\sqrt{2\mathfrak{E}_{\nu}}}\left(\begin{array}[]{c}0\\ \frac{\mathrm{i}\mu B}{\sqrt{\mathfrak{E}_{\nu}+\frac{\Delta m^{2}}{4E_{\nu}}\cos 2\theta}}\\ \sqrt{\mathfrak{E}_{\nu}+\frac{\Delta m^{2}}{4E_{\nu}}\cos 2\theta}\\ -\frac{\frac{\Delta m^{2}}{4E_{\nu}}\sin 2\theta}{\sqrt{\mathfrak{E}_{\nu}+\frac{\Delta m^{2}}{4E_{\nu}}\cos 2\theta}}\end{array}\right),$ (6.21)
are the eigenvectors which are normalized, $|U_{1,2}|^{2}=|V_{1,2}|^{2}=1$.
We suppose that $\nu(0)\equiv\nu_{e}$ initially. Thus, we take $\nu_{0}^{\mathrm{T}}=(1,0,0,0)$ in Eq. (6.2). Using Eq. (6.21), the general solution in Eq. (6.2), satisfying the given initial condition, takes the form,
$\nu(t)=\frac{1}{2\mathfrak{E}_{\nu}}\left(\begin{array}[]{c}\frac{(\mu B)^{2}+\left(\frac{\Delta m^{2}}{4E_{\nu}}\right)^{2}\sin^{2}2\theta}{\mathfrak{E}_{\nu}+\frac{\Delta m^{2}}{4E_{\nu}}\cos 2\theta}e^{-\mathrm{i}\mathfrak{E}_{\nu}t}+\left(\mathfrak{E}_{\nu}+\frac{\Delta m^{2}}{4E_{\nu}}\cos 2\theta\right)e^{\mathrm{i}\mathfrak{E}_{\nu}t}\\ -\mathrm{i}\frac{\Delta m^{2}}{2E_{\nu}}\sin 2\theta\sin\mathfrak{E}_{\nu}t\\ 0\\ -2\mu B\sin\mathfrak{E}_{\nu}t\end{array}\right),$ (6.22)
Based on Eq. (6.22), we get that $\bar{\nu}_{\mu}(t)=-\frac{\mu B}{\mathfrak{E}_{\nu}}\sin\mathfrak{E}_{\nu}t$. The transition probability for the process $\nu_{e}\to\bar{\nu}_{\mu}$ is
$P_{\nu_{e}\to\bar{\nu}_{\mu}}(t)=|\bar{\nu}_{\mu}(t)|^{2}=\frac{(\mu B)^{2}}{(\mu B)^{2}+\left(\frac{\Delta m^{2}}{4E_{\nu}}\right)^{2}}\sin^{2}\left(\sqrt{(\mu B)^{2}+\left(\frac{\Delta m^{2}}{4E_{\nu}}\right)^{2}}t\right).$ (6.23)
It is interesting to mention that the transition probability in Eq. (6.23) does not depend on the vacuum mixing angle.
Now, we can compare the quantum mechanical and quantum field theory approaches to neutrino spin-flavor oscillations. If both neutrinos and charged leptons, which are used to produce and detect neutrinos, are ultrareletivistic, we can put $E_{\nu}=E$, $\mathfrak{E}_{\nu}=\mathfrak{E}$, and $t\approx L$ in Eq. (6.23). Then, we omit the small correction $\delta P_{\nu_{e}\to\bar{\nu}_{\mu}}$ in Eq. (5.17). In this situation, $P_{\nu_{e}\to\bar{\nu}_{\mu}}^{(0)}$ in Eq. (5.16) coincides with the transition probability in Eq. (6.23), i.e. both approaches provide the same description of spin-flavor oscillations.
7 Conclusion
In the present work, we have developed the QFT description of neutrino spin-flavor oscillations in a magnetic field. We were based on the QFT formalism, which was originally developed in Refs. [13, 14, 15], where neutrinos are virtual particles. This method was used in Ref. [13] to study neutrino-to-antineutrino oscillations in vacuum. If one deals with Majorana neutrinos with transition magnetic moments, $\nu\leftrightarrow\bar{\nu}$ oscillations are analogous to spin-flavor transitions in a magnetic field. We have adapted the results of Ref. [13] for the description of spin-flavor oscillations.
In Sec. 2, we have reminded how Majorana neutrinos can interact with electromagnetic fields. Considering the system of two neutrinos, we have obtained the wave equations for neutrino mass eigenstates in an external magnetic field. The magnetic interaction was shown to mix different mass eigenstates, as well as induce particle-to-antiparticle transitions.
The basics of the QFT approach for neutrino spin-flavor oscillations have been developed in Sec. 3. We have obtained the matrix element corresponding to the transition $\nu_{\beta}\to\bar{\nu}_{\alpha}$ in an external magnetic field. The dressed propagators for neutrino mass eigenstates, accounting for the magnetic interaction, have been derived in Sec. 4 for neutrinos with arbitrary energies. Then, we have considered the case of ultrarelativistic particles. Note that we have obtained both diagonal and nondiagonal dressed propagators in Sec. 4.
Finally, we have derived the transition probability for $\nu_{e}\to\bar{\nu}_{\mu}$ oscillations in a magnetic field in Sec. 5 within the QFT approach. This transition probability has the leading term, which was shown in Sec. 6 to coincide with the result of the standard quantum mechanical approach. The correction to the leading term, arising from the QFT description, has been also obtained in Sec. 5; cf. Eq. (5.17). Unfortunately, this correction is vanishing for any oscillations channels and any realistic magnetic fields.
The main goal of the present work is methodological. It is aimed to demonstrate that neutrino spin-flavor oscillations can be described purely with help of the standard QFT methods if reasonable approximation are made. The major assumption is the ultrarelativity of charged leptons which are used to produce and detect virtual neutrinos. The next important assumption, which was also noticed in Ref. [19], is the consideration of two mass eigenstates only. This feature allows one to avoid the branching of the Feynman graphs in Fig. 2.
We also make an interesting observation about neutrino oscillations in external fields. The neutrino interaction with background matter or with a magnetic field can be rather trivially incorporated to the neutrino vacuum oscillations description. Indeed, one just adds a linear term with a certain external field to the vacuum effective Hamiltonian for flavor neutrinos. However, from the point of view of QFT, the effect of external fields on neutrino oscillations is an essentially nonperturbative phenomenon. The correct form of the transition probability is obtained only if the infinite number of terms in Eq. (4) is summed up.
Acknowledgments
I am thankful to M. Deka for the help with the Feynman diagram drawing.
Appendix A Propagators of Weyl neutrinos in vacuum
In this Appendix, we study the propagators of Majorana neutrinos in vacuum. Using the Weyl form of the neutrino wavefunctions, we represent the propagator for left-handed neutrinos and derive the propagator for right-handed antineutrinos.
The propagator of the field $\eta_{a}$, which corresponds to left-handed neutrinos, was obtained in Ref. [25] (see also Ref. [19], where this propagator was derived in the presence of background matter). Using this result, we represent the propagator in question in the form,
$S_{a}^{(\eta)}(x)=\mathrm{i}\int\frac{d^{4}p}{(2\pi)^{4}}e^{-\mathrm{i}px}\frac{1}{2}\left\{\left[1-\frac{(\bm{\sigma}\mathbf{p})}{E_{a}}\right]\frac{1}{p_{0}-E_{a}+\mathrm{i}0}+\left[1+\frac{(\bm{\sigma}\mathbf{p})}{E_{a}}\right]\frac{1}{p_{0}+E_{a}-\mathrm{i}0}\right\},$ (A.1)
[…]
$S_{a\mathrm{L}}(p)=\frac{\mathrm{i}}{2}\left[1-\frac{(\bm{\sigma}\mathbf{p})}{E_{a}}\right]\frac{1}{p_{0}-E_{a}+\mathrm{i}0},$ (A.2)
is the propagator of left-handed neutrinos which is used in Sec. 4.
The the wavefunction $\xi_{a}$, corresponding to a right handed antineutrino, obeys the equation,
$\mathrm{i}\dot{\xi}_{a}-(\bm{\sigma}\mathbf{p})\xi_{a}-\mathrm{i}m_{a}\sigma_{2}\xi_{a}^{*}=0,$ (A.3)
which can be obtained from the Dirac equation for a Majorana neutrino if we take that the bispinor has the form, $\psi_{a}^{\mathrm{T}}=(\xi_{a},-\mathrm{i}\sigma_{2}\xi_{a}^{*})$. Such a bispinor satisfies the Majorana condition, $(\psi_{a})^{c}=\psi_{a}$. The general solution of Eq. (A.3) has the form,
$\xi_{a}(\mathbf{x},t)=$ $\int\frac{\mathrm{d}^{3}p}{(2\pi)^{3/2}}\sqrt{\frac{E_{a}+p}{E_{a}}}\Big{[}\left(a_{a+}w_{+}+A_{a-}a_{a-}w_{-}\right)e^{-\mathrm{i}E_{a}t+\mathrm{i}\mathbf{px}}$
$+\left(a_{a-}^{\dagger}w_{+}+A_{a+}a_{a+}^{\dagger}w_{-}\right)e^{\mathrm{i}E_{a}t-\mathrm{i}\mathbf{px}}\Big{]},$ (A.4)
where
$A_{a\pm}=\pm\frac{m_{a}}{E_{a}+p},$ (A.5)
and
$w_{+}=\left(\begin{array}[]{c}e^{-\mathrm{i}\varphi/2}\cos\vartheta/2\\ e^{\mathrm{i}\varphi/2}\sin\vartheta/2\end{array}\right),\quad w_{-}=\left(\begin{array}[]{c}-e^{-\mathrm{i}\varphi/2}\sin\vartheta/2\\ e^{\mathrm{i}\varphi/2}\cos\vartheta/2\end{array}\right),$ (A.6)
are the helicity amplitudes, which satisfy the relation, $w_{\pm}\otimes w_{\pm}^{\dagger}=\frac{1}{2}\left[1\pm\bm{\sigma}\hat{p}\right]$. In Eq. (A.6), the angles $\vartheta$ and $\varphi$ fix the direction of the neutrino momentum $\mathbf{p}$. The creation and annihilation operators, $a_{a\pm}^{\dagger}(\mathbf{p})$ and $a_{a\pm}(\mathbf{p})$, obey the anticommurators,
$\left\{a_{a\pm}(\mathbf{p}),a_{b\pm}^{\dagger}(\mathbf{q})\right\}_{+}=\delta_{ab}\delta(\mathbf{p}-\mathbf{q}),$ (A.7)
with the rest of the anticommutators being equal to zero.
Then, we define the propagator of $\xi_{a}$ as
$S_{a}^{(\xi)}(x-y)=\theta(x_{0}-y_{0})\left\langle 0|\xi(x)\xi^{\dagger}(y)|0\right\rangle-\theta(y_{0}-x_{0})\left\langle 0|\xi^{*}(y)\xi^{\mathrm{T}}(x)|0\right\rangle.$ (A.8)
Using Eqs. (A)-(A.7) and making calculation analogous to those in Refs. [25, 19], we cast Eq. (A.8) to the form,
$S_{a}^{(\xi)}(x)=\mathrm{i}\int\frac{d^{4}p}{(2\pi)^{4}}e^{-\mathrm{i}px}\frac{1}{2}\left\{\left[1+\frac{(\bm{\sigma}\mathbf{p})}{E_{a}}\right]\frac{1}{p_{0}-E_{a}+\mathrm{i}0}+\left[1-\frac{(\bm{\sigma}\mathbf{p})}{E_{a}}\right]\frac{1}{p_{0}+E_{a}-\mathrm{i}0}\right\}.$ (A.9)
Similarly to $S_{a}^{(\eta)}$ in Eq. (A.1), we decompose $S_{a}^{(\xi)}$ in Eq. (A.9) as $S_{a}^{(\xi)}=S_{a\mathrm{R}}+\tilde{S}_{a\mathrm{L}}$, where
$S_{a\mathrm{R}}(p)=\frac{\mathrm{i}}{2}\left[1+\frac{(\bm{\sigma}\mathbf{p})}{E_{a}}\right]\frac{1}{p_{0}-E_{a}+\mathrm{i}0},$ (A.10)
is the propagator of right-handed antineutrinos which is used in Sec. 4.