### v2

# Quantum field theory treatment of the neutrino spin-flavor precession in a magnetic field

Abstract
We study the spin-flavor precession of neutrinos in a magnetic field within the quantum field theory approach in which neutrinos are virtual particles. Neutrinos are taken to be Majorana particles having a nonzero transition magnetic moment. We derive the dressed propagators of the neutrino mass eigenstates exactly accounting for the magnetic field contribution. The matrix element and the transition probability for the spin-flavor precession are obtained in the approximation of the forwardly scattered charged leptons. The leading term in the transition probability is shown to coincide with the result of the standard quantum mechanical description of neutrino oscillations. We also discuss the quantum field theory contributions to the neutrino dressed propagators and demonstrate that these contributions result in a small correction to the transition probability. The case of charged leptons with arbitrary energies is considered.
1 Introduction
The interaction of neutrinos with external fields can significantly modify the dynamics of neutrino oscillations. From the point of view of the phenomenology, the most important external field, contributing to flavor oscillations, is the electroweak interaction of neutrinos with background matter. It can result in the sizable amplification of the transition probability, known as the Mikheyev–Smirnov–Wolfenstein (MSW) effect [1, 2], that is the most plausible explanation of the solar neutrino deficit [3].
Neutrinos are supposed to be electrically neutral particles. Nevertheless, these particles can interact with electromagnetic fields owing to nonzero magnetic moments which have a purely anomalous origin (see, e.g., Ref. [4]). It should be noted that the electromagnetic properties of Dirac and Majorana neutrinos are completely different. Dirac neutrinos can have both diagonal and transition magnetic moments, whereas Majorana neutrinos can possess only transition ones. Despite the issue of the neutrino nature is still open, neutrinos are likely to be Majorana particles. At least, the smallness of the active neutrino masses is explained more naturally for Majorana neutrinos [5]. The recent advances in the studies of neutrinos electromagnetic properties are reviewed in Ref. [6].
If neutrinos with transition magnetic moments interact with an external magnetic field, they undergo both flavor change and the spin flip. This process is called the neutrino spin-flavor precession [7]. The neutrino spin-flavor precession was studied, e.g., in Ref. [8] to explain the possible anticorrelation between the solar activity and the solar neutrinos flux. The spin-flavor precession accounting for the neutrino interaction with background matter was shown in Ref. [9] to exhibit a resonant behavior analogous to the MSW effect. The application of the predicted effect to solar neutrinos was also discussed in Ref. [9]. Later, the spin-flavor precession was understood in Ref. [10] to play a subdominant role for the solar neutrinos deficit problem. The description of the spin-flavor precession of Majorana neutrinos expressed in terms of the Weyl spinors was carried out in Ref. [11].
As a rule, neutrino oscillations in vacuum and in external fields are described within the quantum mechanical approach, which gives satisfactory results in the majority of situations. Nonetheless, this formalism has some shortcomings which were outlined in Ref. [12]. Numerous attempts to construct the description of neutrino flavor oscillations based on the quantum field theory (QFT) methods are reviewed in Ref. [13].
We mention one of the QFT approaches which was put forward independently in Refs. [14, 15, 16]. Neutrinos are supposed to be virtual particles in this formalism. We also mention Refs. [17, 18], where QFT was applied to study neutrino flavor oscillations in external fields.
The approach in Refs. [14, 15, 16] allows one to treat neutrino oscillations in vacuum since it involves the propagators of the neutrino mass eigenstates. In vacuum, these propagators are diagonal in the neutrino types. If a neutrino interacts with an external field, the mass eigenstates propagators become nondiagonal since external fields mix different mass eigenstates as a rule. Finding such nondiagonal propagators in not a trivial task. A particular case of the spin-flavor precession of Dirac neutrinos, possessing diagonal magnetic moments in the mass basis, was analyzed in frames of QFT in Ref. [19].
Recently, in Ref. [20], we developed a method which enables one to find the propagators of neutrino mass eigenstates in an external field. This approach is based on solving of the Dyson equations for propagators accounting for the interaction with external fields. Using this method in Ref. [20], we described neutrino flavor oscillations in background matter within QFT.
In the present work, we apply the results of Ref. [20] for the studies of the neutrino spin-flavor precession in a magnetic field. Moreover, we consider the situation when the interaction with an external field is essentially nondiagonal in the mass basis. We shall see shortly in Sec. 2 that it is implemented for Majorana neutrinos. Therefore, the application of the results of Ref. [20] is crucial in this case. Note that the alternative formalism, based on the relativistic quantum mechanics, for the description of neutrino flavor oscillations and the spin-flavor precession in various external fields was considered in Ref. [21].
This work is organized in the following way. We remind briefly some basic issues of the Majorana neutrinos electrodynamics in Sec. 2. Then, in Sec. 3, we outline the QFT approach and adapt it for the neutrino spin-flavor precession. We derive the dressed propagators for the neutrino mass eigenstates in a magnetic field by solving the corresponding Dyson equations in Sec. 4. The matrix element and the transition probability for the neutrino spin-flavor precession are computed in Sec. 5. The results of the QFT and the standard quantum mechanical approaches are compared in Sec. 6. Finally, we discuss our results in Sec. 7. The propagators of massive Weyl neutrinos in vacuum are derived in Appendix.
2 Electrodynamics of Majorana neutrinos
In this section, we remind how massive Majorana neutrinos interact with electromagnetic fields.
Let us consider the system of two active flavor neutrinos $\nu_{\lambda}=(\nu_{e},\nu_{\mu})$. The mass matrix of these neutrinos is supposed to be nondiagonal. For simplicity, we take that no sterile neutrinos are present. To diagonalize the mass matrix we introduce the mass eigenstates $\psi_{a}$, $a=1,2$, by means of the matrix transformation,
$\nu_{\lambda}=\sum_{a}U_{\lambda a}\psi_{a},\quad(U_{\lambda a})=\left(\begin{array}[]{cc}\cos\theta&\sin\theta\\ -\sin\theta&\cos\theta\end{array}\right),$ (2.1)
where $\theta$ is the vacuum mixing angle. We assume that $\psi_{a}$ are Majorana particles having masses $m_{a}$.
We suppose that $\psi_{a}$ have the nonzero matrix of magnetic moments $(\mu_{ab})$. This matrix was found in Ref. [22] to be hermitian and purely imaginary. Hence, if we deal with two mass eigenstates, we can represent this matrix as $\mu_{ab}=\mu(\sigma_{2})_{ab}$, where $a,b=1,2$, $\sigma_{2}$ is the Pauli matrix, and $\mu$ is the transition magnetic moment. Then, we suppose that these neutrinos interact with an external electromagnetic field $F_{\mu\nu}=(\mathbf{E},\mathbf{B})$. The wave equations for the mass eigenstates can be formally obtained from the Dirac equation for $\psi_{a}$ [23],
$(\mathrm{i}\gamma^{\mu}\partial_{\mu}-m_{a})\psi_{a}-\frac{\mu_{ab}}{2}\sigma_{\mu\nu}F^{\mu\nu}\psi_{b}=0,$ (2.2)
where $\gamma^{\mu}$, $\gamma^{5}$, and $\sigma_{\mu\nu}=(\mathrm{i}/2)(\gamma_{\mu}\gamma_{\nu}-\gamma_{\nu}\gamma_{\mu})$ are the Dirac matrices. Since the mass eigenstates are taken to be Majorana particles, we represent them in terms of the Weyl spinors $\eta_{a}$, $\psi_{a}^{\mathrm{T}}=(\mathrm{i}\sigma_{2}\eta_{a}^{*},\eta_{a})$. Then, assuming that the electric field is absent and using Eq. (2.2), we derive the wave equations for the mass eigenstates,
$\mathrm{i}\dot{\eta}_{1}+(\bm{\sigma}\mathbf{p})\eta_{1}+\mathrm{i}m_{1}\sigma_{2}\eta_{1}^{*}-\mu(\bm{\sigma}\mathbf{B})\sigma_{2}\eta_{2}^{*}$ $=0,$
$\mathrm{i}\dot{\eta}_{2}+(\bm{\sigma}\mathbf{p})\eta_{2}+\mathrm{i}m_{2}\sigma_{2}\eta_{2}^{*}+\mu(\bm{\sigma}\mathbf{B})\sigma_{2}\eta_{1}^{*}$ $=0,$ (2.3)
where $\mathbf{p}=-\mathrm{i}\nabla$ is the momentum operator and $\bm{\sigma}=(\sigma_{1},\sigma_{2},\sigma_{3})$ are the Pauli matrices.
One can see in Eq. (2) that the neutrino magnetic interaction causes transitions between different mass eigenstates, $1\leftrightarrow 2$. Moreover, the magnetic interaction has the structure similar to the mass terms, which are known to induce neutrino-to-antineutrino oscillations [14, 24]. Indeed, both terms are $\propto\sigma_{2}\eta_{a,b}^{*}$. Thus, neutrinos undergo transitions $1\leftrightarrow\bar{2}$ and $\bar{1}\leftrightarrow 2$, where the bar denotes an antiparticle state, in an external magnetic field.
3 Quantum field theory formalism for neutrino oscillations
In this section, we briefly outline the quantum field theory based formalism for neutrino oscillations in which neutrinos are virtual particles [14, 15, 16]. We adapt this method for the description of the spin-flavor precession.
Following Ref. [14], we assume that the source and the detector of neutrinos, separated by the distance $\mathbf{L}$, consist of heavy nuclei $N$ and $N^{\prime}$. The production and detection of neutrinos is characterized by the interaction of these nuclei with charged leptons $l_{\alpha}^{-}$ and antileptons $l_{\beta}^{+}$, where $\alpha,\beta=e,\mu,\dotsc$. Thus, neutrino oscillations correspond to the process $l_{\beta}^{-}+N\to\text{neutrinos}\to l_{\alpha}^{+}+\tilde{N}^{\prime}$. If the flavor content of charged leptons changes, i.e. $\beta\neq\alpha$, one can speak about neutrino flavor oscillations. If electric charges of leptons also change, e.g., we have an electron in a source and antimuon in a detector, such a process is allowed in vacuum for Majorana neutrinos. It was connected to neutrino-to-antineutrino oscillations in Refs. [14, 24]. However, if we study the transition $l_{\beta}^{-}+N\to\text{neutrinos}\to l_{\alpha}^{+}+\tilde{N}^{\prime}$ in an external field which neutrinos interact with, it is more appropriate to call it the neutrino spin-flavor precession. The considered process is schematically depicted in Fig. 1.
Figure 1: The schematic Feynman diagram for the reaction $l_{\beta}^{-}+N\to\text{neutrinos}\to l_{\alpha}^{+}+\tilde{N}^{\prime}$. The neutrino line has two counter-directional arrows since we deal with the Majorana particles. Moreover, the neutrino line is broad since the mass eigenstates interact with a magnetic field [see Eq. (2)]. Note that we take into account the magnetic interaction exactly (see Sec. 4).
The $S$-matrix element of the reaction in question reads [14]
$S=-\frac{1}{2}\left(\sqrt{2}G_{\mathrm{int}}\right)^{2}\int\mathrm{d}^{4}x\mathrm{d}^{4}y\Big{\langle}l_{\alpha}^{+},\tilde{N},\tilde{N}^{\prime}\Big{|}T\left\{j^{\mu}(x)J_{\mu}(x)j^{\nu}(y)J_{\nu}(y)\right\}\Big{|}l_{\beta}^{-},N,N^{\prime}\Big{\rangle},$ (3.1)
where $G_{\mathrm{int}}$ is the coupling constant,
$j_{\mu}=\sum_{\lambda}\bar{\nu}_{\lambda}\gamma_{\mu}^{\mathrm{L}}l_{\lambda},$ (3.2)
is the leptons current, $\gamma_{\mu}^{\mathrm{L}}=\gamma_{\mu}(1-\gamma^{5})/2$, and $J_{\mu}(x)$ is the nuclear current operator. Then, we represent the operator valued wave function of leptons as
$l_{\lambda}(x)=\int\frac{\mathrm{d}^{3}p}{(2\pi)^{3/2}}\left[a_{\lambda}(p)u_{\lambda}(p)e^{-\mathrm{i}p_{\lambda}x}+b_{\lambda}^{\dagger}(p)v_{\lambda}(p)e^{\mathrm{i}p_{\lambda}x}\right],$ (3.3)
where $b_{\lambda}^{\dagger}(p)$ and $a_{\lambda}(p)$ are the creation and annihilation operators of antileptons and leptons, $u_{\lambda}(p)$ and $v_{\lambda}(p)$ are the basis spinors, $p_{\lambda}^{\mu}=(E_{\lambda},\mathbf{p})$, $E_{\lambda}=\sqrt{\mathfrak{m}_{\lambda}^{2}+p^{2}}$ is the lepton energy, and $\mathfrak{m}_{\lambda}$ is its mass. We take that $\left|l_{\beta}^{-}\right\rangle=a_{\beta}^{\dagger}(p_{\beta})\left|0\right\rangle$ and $\left\langle l_{\alpha}^{+}\right|=\left\langle 0\right|b_{\alpha}(p_{\alpha})$, i.e. we deal with the plane waves incoming and outgoing leptons. Eventually, we rewrite Eq. (3.1) in the following way:
$S=$ $2G_{\mathrm{int}}^{2}\sum_{ab}U_{\alpha a}^{*}U_{\beta b}^{*}v_{\alpha}^{\mathrm{T}}(p_{\alpha})(\gamma_{\mathrm{L}}^{\mu})^{\mathrm{T}}$
$\times\left(\int\mathrm{d}^{4}x\mathrm{d}^{4}ye^{\mathrm{i}p_{\alpha}x-\mathrm{i}p_{\beta}y}J_{\mu}(x)J_{\nu}(y)\big{\langle}0\big{|}T\left\{(\bar{\psi}_{a}(x))^{\mathrm{T}}\bar{\psi}_{b}(y)\right\}\big{|}0\big{\rangle}\right)\gamma_{\mathrm{L}}^{\nu}u_{\beta}(p_{\beta}),$ (3.4)
where we account for Eqs. (3.2) and (3.3). In Eq. (3), we use the mean values of the nuclear currents $J_{\mu}(x)=\langle\tilde{N}^{\prime}|J_{\mu}|N^{\prime}\rangle$ and $J_{\nu}(y)=\langle\tilde{N}|J_{\nu}|N\rangle$.
We assume that the nuclei in the source and the detector are quite heavy. Hence, their positions, $\mathbf{x}_{1}$ and $\mathbf{x}_{2}$, are unchanged, with $\mathbf{x}_{2}-\mathbf{x}_{1}=\mathbf{L}$. In this case, the nuclear currents in Eq. (3) have the form,
$J_{\mu}(x_{0},\mathbf{x})\propto\delta_{\mu 0}\delta(\mathbf{x}-\mathbf{x}_{2}),\quad J_{\nu}(y_{0},\mathbf{y})\propto\delta_{\nu 0}\delta(\mathbf{y}-\mathbf{x}_{1}).$ (3.5)
We take that the neutrino mass eigenstates are Majorana particles, i.e. $\left(\psi_{a}\right)^{c}=C\bar{\psi}_{a}^{\mathrm{T}}=\psi_{a}$, where $C$ is the charge conjugation matrix. In this case,
$v_{\alpha}^{\mathrm{T}}(p_{\alpha})(\gamma_{\mathrm{L}}^{0})^{\mathrm{T}}\left\langle 0\left|T\left\{(\bar{\psi}_{a}(x))^{\mathrm{T}}\bar{\psi}_{b}(y)\right\}\right|0\right\rangle\gamma_{\mathrm{L}}^{0}u_{\beta}(p_{\beta})\\ =\kappa_{\bar{\alpha}\mathrm{R}}^{\dagger}(p_{\alpha})\left\langle 0\left|T\left\{\xi_{a}(x)\eta_{b}^{\dagger}(y)\right\}\right|0\right\rangle\kappa_{\beta\mathrm{L}}(p_{\beta}).$ (3.6)
Here we use the representations
$\psi_{a}=\left(\begin{array}[]{c}\xi_{a}\\ \eta_{a}\end{array}\right),\quad u_{\alpha,\beta}(p_{\alpha,\beta})=\left(\begin{array}[]{c}\kappa_{\bar{\alpha},\beta\mathrm{R}}(p_{\alpha,\beta})\\ \kappa_{\bar{\alpha},\beta\mathrm{L}}(p_{\alpha,\beta})\end{array}\right).$ (3.7)
Note that we keep $\xi_{a}$ instead of $\mathrm{i}\sigma_{2}\eta_{a}^{*}$ in Eq. (3.7). We also use the known expressions, $\gamma_{\mathrm{R}}^{\mu}=C(\gamma_{\mathrm{L}}^{\mu})^{\mathrm{T}}C$ and $v_{\alpha}(p_{\alpha})=\left(u_{\alpha}(p_{\alpha})\right)^{c}$, to derive Eq. (3.6).
Using Eq. (3.6) and performing the integrations as in Ref. [14], we represent Eq. (3) in the form,
$S=4\pi G_{\mathrm{int}}^{2}e^{-i\mathbf{p}_{\alpha}\mathbf{x}_{2}+i\mathbf{p}_{\beta}\mathbf{x}_{1}}\delta(E_{\alpha}-E_{\beta})\mathcal{M}_{\beta\to\bar{\alpha}},$ (3.8)
where the matrix element is
$\mathcal{M}_{\beta\to\bar{\alpha}}=\sum_{ab}U_{\alpha a}^{*}U_{\beta b}^{*}\kappa_{\bar{\alpha}\mathrm{R}}^{\dagger}(p_{\alpha})\left(\int\frac{\mathrm{d}^{3}q}{(2\pi)^{3}}\Sigma_{ab}(E,\mathbf{q})e^{\mathrm{i}\mathbf{qL}}\right)\kappa_{\beta\mathrm{L}}(p_{\beta}).$ (3.9)
Here, $E=(E_{\alpha}+E_{\beta})/2$ and $\Sigma_{ab}(q_{0},\mathbf{q})$ are the 4D Fourier images of the nondiagonal propagators,
$\Sigma_{ab}(x-y)=\left\langle 0\left|T\left\{\xi_{a}(x)\eta_{b}^{\dagger}(y)\right\}\right|0\right\rangle.$ (3.10)
If we consider utrarelativistic incoming and outgoing leptons, then, in Eq. (3.9), we should replace $\kappa_{\beta\mathrm{L}}$ and $\kappa_{\bar{\alpha}\mathrm{R}}$ with the helicity states $\kappa_{\beta-}$ and $\kappa_{\bar{\alpha}+}$.
The probability of the neutrino spin-flavor precession is $P_{\beta\to\bar{\alpha}}\propto|\mathcal{M}_{\beta\to\bar{\alpha}}|^{2}$. In case of the Majorana neutrinos motion in vacuum, the propagators are diagonal, $\Sigma_{ab}\propto\delta_{ab}$. In this situation, the process $\beta\to\bar{\alpha}$ corresponds to neutrino-to-antineutrino oscillations, as we mentioned earlier. The probability of $\nu_{\alpha}\to\bar{\nu}_{\beta}$ oscillations was found in Refs. [14, 24] to be of the order $\sim(m_{1,2}/E)^{2}$.
If Majorana neutrinos have a transition magnetic moment, as one can see in Eq. (2), the wave equations for $\eta_{1}$ and $\eta_{2}$ are coupled by the magnetic interaction. This interaction induces the transitions between $\eta_{1,2}$ and $\mathrm{i}\sigma_{2}\eta_{2,1}^{*}=\xi_{2,1}$. It means that both diagonal, $\Sigma_{11}$ and $\Sigma_{22}$, and nondiagonal, $\Sigma_{12}$ and $\Sigma_{21}$, propagators in Eq. (3.10) can contribute to the transition probability. In case we study the spin-flavor precession like $\nu_{e}\to\bar{\nu}_{\mu}$, the matrix element in Eq. (3.9) takes the form,
$\mathcal{M}_{e\to\bar{\mu}}=\kappa_{\bar{\mu}\mathrm{R}}^{\dagger}(p_{\mu})\int\frac{\mathrm{d}^{3}q}{(2\pi)^{3}}e^{\mathrm{i}\mathbf{qL}}\left[\frac{1}{2}\left(\Sigma_{22}-\Sigma_{11}\right)\sin 2\theta+\cos^{2}\theta\Sigma_{21}-\sin^{2}\theta\Sigma_{12}\right]\kappa_{e\mathrm{L}}(p_{e}),$ (3.11)
where we use Eq. (2.1). Our main goal is to find the propagators $\Sigma_{ab}$ in Eq. (3.11) in the presence of the magnetic field.
4 Propagators of massive Majorana neutrinos in a magnetic field
In this section, we derive the dresses propagators of Majorana neutrinos interacting with a magnetic field. For this purpose, we use the method developed in Ref. [20], where the propagators in question are the solutions of the Dyson equations exactly accounting for the interaction with an external field. First, we derive the Dyson equations within the perturbative approach, with all the terms in the perturbation series being summed up. Then, the Dyson equations are solved exactly for ultrarelativistic neutrinos. We deal with two mass eigenstates since we mentioned in Ref. [20] that extension of the formalism for the greater number of neutrinos is ambiguous.
The magnetic interaction in Eq. (2) mixes both opposite helicities and different mass eigenstates. Thus, if we have a left-handed lepton in a source and a right-handed one in a detector, like in Eq. (3.11), the dressed propagators are the sums of the following perturbative series:
$-\mathrm{i}\Sigma_{12}=$ $(-\mathrm{i}S_{1\mathrm{R}})V(-\mathrm{i}S_{2\mathrm{L}})+(-\mathrm{i}S_{1\mathrm{R}})V(-\mathrm{i}S_{2\mathrm{L}})(-V)(-\mathrm{i}S_{1\mathrm{R}})V(-\mathrm{i}S_{2\mathrm{L}})+\dotsb,$
$-\mathrm{i}\Sigma_{21}=$ $(-\mathrm{i}S_{2\mathrm{R}})(-V)(-\mathrm{i}S_{1\mathrm{L}})+(-\mathrm{i}S_{2\mathrm{R}})(-V)(-\mathrm{i}S_{1\mathrm{L}})V(-\mathrm{i}S_{2\mathrm{R}})(-V)(-\mathrm{i}S_{1\mathrm{L}})+\dotsb,$
$-\mathrm{i}\Sigma_{11}=$ $-\mathrm{i}S_{1\mathrm{L}}+(-\mathrm{i}S_{1\mathrm{L}})V(-\mathrm{i}S_{2\mathrm{R}})(-V)(-\mathrm{i}S_{1\mathrm{L}})+\dotsb,$
$-\mathrm{i}\Sigma_{22}=$ $-\mathrm{i}S_{2\mathrm{L}}+(-\mathrm{i}S_{2\mathrm{L}})(-V)(-\mathrm{i}S_{1\mathrm{R}})V(-\mathrm{i}S_{2\mathrm{L}})+\dotsb,$ (4.1)
where $V=\mathrm{i}\mu(\bm{\sigma}\mathbf{B})$ is the magnetic interaction and $S_{a\mathrm{L,R}}$ are the propagators of left-handed neutrinos and right-handed antineutrinos. Since the magnetic interaction contributes to the transitions $a\leftrightarrow\bar{b}$ only, the diagonal propagators $S_{a\mathrm{L,R}}$ should be taken in vacuum. We give them in Eqs. (A.2) and (A.10). We also take into account in Eq. (4) that the magnetic interaction contributes to the wave equations for $\eta_{1}$ and $\eta_{2}$ with opposite signs; cf. Eq. (2). The perturbative series in Eq. (4) are illustrated in Fig. 2.
Figure 2: Feynman diagrams corresponding to Eq. (4). 4D Fourier images of the bare propagators $\mathrm{i}S_{a\mathrm{R,L}}$, given in Eqs. (A.2) and (A.10), are depicted by thin solid lines [25, pp. 468–472]. The broad solid lines correspond to the dressed propagators $\mathrm{i}\Sigma_{ab}$. The potential of the magnetic interaction is $V=\mathrm{i}\mu(\bm{\sigma}\mathbf{B})$.
The series in Eq. (4) can be summed up, giving the result
$\mathrm{i}\Sigma_{12}^{-1}$ $=-(S_{1\mathrm{R}}VS_{2\mathrm{L}})^{-1}+V,\quad\mathrm{i}\Sigma_{21}^{-1}=(S_{2\mathrm{R}}VS_{1\mathrm{L}})^{-1}-V,$
$\Sigma_{11}^{-1}$ $=S_{1\mathrm{L}}^{-1}-VS_{2\mathrm{R}}V,\quad\Sigma_{22}^{-1}=S_{2\mathrm{L}}^{-1}-VS_{1\mathrm{R}}V.$ (4.2)
We notice the diagonal propagators in Eq. (4) do not contribute to the matrix element since $\Sigma_{aa}$ preserves the neutrino helicity, whereas $\mathcal{M}_{e\to\bar{\mu}}$ implies the spin-flip. We provide the detailed calculation for $\Sigma_{12}$ and $\Sigma_{21}$ only.
[…]
We provide the detailed calculations of $\Sigma_{12}$ only. Using Eqs. (A.2) and (A.10), the exact solution of the Dyson equation for $\Sigma_{12}$ in Eq. (4) has the form,
$\Sigma_{12}(E,\mathbf{q})=$ $-\frac{1}{4\alpha\varsigma_{1}\varsigma_{2}\mu^{2}(\hat{q}\mathbf{B})^{2}+(\mu B)^{2}[\alpha^{2}(1-\varsigma_{1}^{2})(1-\varsigma_{2}^{2})-2\alpha(1+\varsigma_{1}\varsigma_{2})+1]}$
$\times\big{\{}\alpha(\varsigma_{1}-\varsigma_{2})\mu(\hat{q}\mathbf{B})+(\bm{\sigma}\cdot[\mathrm{i}\alpha(\varsigma_{1}+\varsigma_{2})\mu(\hat{q}\times\mathbf{B})-2\alpha\varsigma_{1}\varsigma_{2}\mu(\hat{q}\mathbf{B})\hat{q}$
$+\mu\mathbf{B}(\alpha(1+\varsigma_{1}\varsigma_{2})-1)])\big{\}},$ (4.3)
where
$\alpha=\frac{4(E-E_{1}+\mathrm{i}0)(E-E_{2}+\mathrm{i}0)}{\mu^{2}B^{2}(1-\varsigma_{1}^{2})(1-\varsigma_{2}^{2})},$ (4.4)
and $E_{1,2}=\sqrt{q^{2}+m_{1,2}^{2}}$. For ultrarelativistic neutrinos, one has that $\varsigma_{1,2}\to 1$, and Eq. (4) takes the form,
$\Sigma_{12}=-\frac{\mu(\bm{\sigma}\cdot[\mathbf{B}+\mathrm{i}(\hat{q}\times\mathbf{B})-(\hat{q}\mathbf{B})\hat{q}])}{2[(E-E_{1}+\mathrm{i}0)(E-E_{2}+\mathrm{i}0)-\mu^{2}B^{2}+\mu^{2}(\hat{q}\mathbf{B})^{2}]}.$ (4.5)
Analogously to Eq. (4.5), one gets that $\Sigma_{21}=-\Sigma_{12}$.
5 Transition probability of the spin-flavor precession
In this section, we calculate the 3D Fourier transform of the dressed propagators, derived in Sec. 4, which the matrix element depends on; cf. Eq. (3.11). Then, we compute the transition probability for the $\nu_{e}\to\bar{\nu}_{\mu}$ precession.
We present the detailed calculation of the $\Sigma_{12}$ contribution to the matrix element. The magnetic field is taken to be transverse with respect $\mathbf{L}$, $\mathbf{B}\perp\mathbf{L}$. For instance, we suppose that $\mathbf{L}=L\mathbf{e}_{z}$ and $\mathbf{B}=B\mathbf{e}_{x}$, where $\mathbf{e}_{x,z}$ are the unit vectors along the $x$- and $z$-axes. In the quantum mechanical approach, a neutrino is on the mass shell, i.e. the neutrino momentum is also perpendicular to $\mathbf{B}$. In the QFT treatment, a neutrino is a virtual particle. Thus, $\mathbf{q}$ is arbitrary and, hence, $(\hat{q}\cdot\mathbf{B})\neq 0$ in Eq. (4.5).
We should keep the terms (i) $\propto\mu(\bm{\sigma}\cdot\hat{q})(\hat{q}\cdot\mathbf{B})$ and (ii) $\propto\mu^{2}(\hat{q}\cdot\mathbf{B})^{2}$ in the numerator and the denominator of $\Sigma_{12}$ in Eq. (4.5) since this kind of terms is of the same order of magnitude as other terms in the propagator. At least, one has no reasons to neglect the terms (i) and (ii) before the calculation of the matrix element.
To evaluate the contribution of the terms (i) and (ii) to the matrix element we adopt the forward scattering approximation, i.e. we assume that $\mathbf{p}_{e}\upuparrows\mathbf{p}_{\bar{\mu}}\upuparrows\mathbf{L}$ to simplify some integrals computation. This approximation is valid for ultrarelativistic charged leptons. The leptons with arbitrary energies are considered shortly in Sec. 5.1.
Following Ref. [20], we use the cylindrical coordinates $(\rho,\phi,z)$ to compute the integral in Eq. (3.11), $\mathbf{q}=\rho(\cos\phi\mathbf{e}_{x}+\sin\phi\mathbf{e}_{y})+z\mathbf{e}_{z}$. In the forward scattering approximation, the spinors of the charged leptons are $\kappa_{e\mathrm{L}}^{\mathrm{T}}=(0,1)$ and $\kappa_{\bar{\mu}\mathrm{R}}^{\mathrm{T}}=(1,0)$. After averaging the contribution of $\Sigma_{12}$ to the matrix element over $\kappa_{\bar{\mu}\mathrm{R}}$ and $\kappa_{e\mathrm{L}}$, we can drop some terms which vanish in the integration over $\phi$ because of the symmetry properties of the integrand. Therefore, the remainung nontrivial contribution reads
$\kappa_{\bar{\mu}\mathrm{R}}^{\dagger}\Sigma_{12}\kappa_{e\mathrm{L}}=\frac{f}{\mathcal{E}_{+}(\mathbf{q})-\mathcal{E}_{-}(\mathbf{q})}\left(\frac{1}{E-\mathcal{E}_{+}(\mathbf{q})+\mathrm{i}0}-\frac{1}{E-\mathcal{E}_{-}(\mathbf{q})+\mathrm{i}0}\right),$ (5.1)
where
$f(\rho,\phi,z)=-\frac{\mu B}{2}\left(1+\frac{z}{\sqrt{\rho^{2}+z^{2}}}-\frac{\rho^{2}\cos^{2}\phi}{\rho^{2}+z^{2}}\right),$ (5.2)
and
$\mathcal{E}_{\pm}(\mathbf{q})=\bar{E}\pm\sqrt{\Delta E^{2}+(\mu B)^{2}\left(1-\frac{\rho^{2}\cos^{2}\phi}{\rho^{2}+z^{2}}\right)}.$ (5.3)
Here $\bar{E}=[E_{2}(\mathbf{q})+E_{1}(\mathbf{q})]/2$ and $\Delta E=[E_{2}(\mathbf{q})-E_{1}(\mathbf{q})]/2$.
Let us compute one of the integrals, which contributes to Eq. (5.1). Namely
$I_{+}=$ $\frac{1}{(2\pi)^{3}}\int_{0}^{2\pi}\mathrm{d}\phi\int_{0}^{\infty}\rho\mathrm{d}\rho\int_{-\infty}^{+\infty}\mathrm{d}z\frac{fe^{\mathrm{i}zL}}{(\mathcal{E}_{+}-\mathcal{E}_{-})(E-\mathcal{E}_{+}+\mathrm{i}0)}$
$=-\frac{\mathrm{i}}{(2\pi)^{2}}\int_{0}^{2\pi}\mathrm{d}\phi\sum_{\left\{z_{+}\right\}}\int_{0}^{\infty}\rho\mathrm{d}\rho\left.\frac{f(\rho,\phi,z_{+})e^{\mathrm{i}z_{+}L}}{(\mathcal{E}_{+}-\mathcal{E}_{-})\frac{\mathrm{d}\mathcal{E}_{+}}{\mathrm{d}z}}\right|_{z=z_{+}},$ (5.4)
where $z_{+}(\rho,\phi)$ are the roots of the equation $E-\mathcal{E}_{+}(\rho,\phi,z)+\mathrm{i}0=0$. These roots are in the upper half-plane. Since we take that $L>0$, we close the contour in the upper half-plane while integrating over $z$.
We define the new quantity
$\mathcal{E}_{+}(E)=E+\sqrt{\left(\frac{\Delta m^{2}}{4E}\right)^{2}+(\mu B)^{2}\left(1-\frac{\rho^{2}}{E^{2}}\cos^{2}\phi\right)},$ (5.5)
where $\Delta m^{2}=m_{2}^{2}-m_{1}^{2}>0$. Note that $\mathcal{E}_{+}(E)\neq\mathcal{E}_{+}(\mathbf{q})$; cf. Eq. (5.3). We can approximately set that $\left.\frac{\mathrm{d}\mathcal{E}_{+}(\mathbf{q})}{\mathrm{d}z}\right|{}_{z=z_{+}}\approx\frac{z_{+}}{\mathcal{E}_{+}(E)}$ and $\mathcal{E}_{+}^{2}(E)\approx\rho^{2}+z_{+}^{2}$. Moreover, for ultrarelativistic neutrinos, having $\mathfrak{E}\ll E$, and a rather weak magnetic field, obeying $(\mu B)^{2}\ll E\mathfrak{E}$, one has that
$\mathcal{E}_{+}(E)\approx E+\mathfrak{E}\left(1-\frac{(\mu B)^{2}\rho^{2}\cos^{2}\phi}{2E^{2}\mathfrak{E}^{2}}\right),$ (5.6)
where
$\mathfrak{E}=\sqrt{(\mu B)^{2}+\left(\frac{\Delta m^{2}}{4E}\right)^{2}}.$ (5.7)
Using Eq. (5.6), one gets the following expressions for the poles in the explicit form:
$z_{+}$ $=\sqrt{1+\frac{(\mu B)^{2}}{E\mathfrak{E}}\cos^{2}\phi}\times\begin{cases}\sqrt{\rho_{0}^{2}-\rho^{2}},&\text{if}\ \rho<\rho_{0},\\ \mathrm{i}\sqrt{\rho^{2}-\rho_{0}^{2}},&\text{if}\ \rho>\rho_{0},\end{cases}$ (5.8)
where $\rho_{0}=(E+\mathfrak{E})/\sqrt{1+\frac{(\mu B)^{2}}{E\mathfrak{E}}\cos^{2}\phi}$.
Using Eqs. (5.2) and (5.8), we rewrite Eq. (5) as
$I_{+}=$ $\frac{\mathrm{i}\mu B(E+\mathfrak{E})}{16\pi^{2}\mathfrak{E}}\int_{0}^{2\pi}\frac{\mathrm{d}\phi}{\sqrt{1+\frac{(\mu B)^{2}}{E\mathfrak{E}}\cos^{2}\phi}}\Bigg{[}\int_{0}^{\rho_{0}}\frac{\rho\mathrm{d}\rho\exp\left(\mathrm{i}\tilde{L}\sqrt{\rho_{0}^{2}-\rho^{2}}\right)}{\sqrt{\rho_{0}^{2}-\rho^{2}}}$
$\times\left(1-\rho^{2}\cos^{2}\phi\left(\frac{1}{(E+\mathfrak{E})^{2}}-\frac{(\mu B)^{2}}{2E^{2}\mathfrak{E}^{2}}\right)+\frac{\sqrt{\rho_{0}^{2}-\rho^{2}}}{\tilde{E}\left(1+\frac{\mathfrak{E}}{E}\right)}\right)$
$-\mathrm{i}\int_{\rho_{0}}^{\infty}\frac{\rho\mathrm{d}\rho\exp\left(-\tilde{L}\sqrt{\rho^{2}-\rho_{0}^{2}}\right)}{\sqrt{\rho^{2}-\rho_{0}^{2}}}$
$\times\left(1-\rho^{2}\cos^{2}\phi\left(\frac{1}{(E+\mathfrak{E})^{2}}-\frac{(\mu B)^{2}}{2E^{2}\mathfrak{E}^{2}}\right)+\frac{\mathrm{i}\sqrt{\rho^{2}-\rho_{0}^{2}}}{\tilde{E}\left(1+\frac{\mathfrak{E}}{E}\right)}\right)\Bigg{]},$ (5.9)
where
$\tilde{L}=L\sqrt{1+\frac{(\mu B)^{2}}{E\mathfrak{E}}\cos^{2}\phi},\quad\tilde{E}=\frac{E}{\sqrt{1+\frac{(\mu B)^{2}}{E\mathfrak{E}}\cos^{2}\phi}}.$ (5.10)
Making the integrations over $\rho$, we represent Eq. (5) in the form,
$I_{+}\approx$ $\frac{\mu B(E+\mathfrak{E})e^{\mathrm{i}(E+\mathfrak{E})}}{8\pi{}^{2}\mathfrak{E}L}\int_{0}^{2\pi}\frac{\mathrm{d}\phi}{1+\frac{(\mu B)^{2}}{E\mathfrak{E}}\cos^{2}\phi}$
$\approx\frac{\mu B(E+\mathfrak{E})e^{\mathrm{i}(E+\mathfrak{E})L}}{4\pi\mathfrak{E}L}\left(1-\frac{(\mu B)^{2}}{2E\mathfrak{E}}\right),$ (5.11)
where we assume that the distance between the source and the detector is great, $L\gg E^{-1}$.
Analogously to Eq. (5), one computes the integral,
$I_{-}=\int\frac{\mathrm{d}^{3}q}{(2\pi)^{3}}\frac{fe^{\mathrm{i}\mathbf{qL}}}{(\mathcal{E}_{+}-\mathcal{E}_{-})(E-\mathcal{E}_{-}+\mathrm{i}0)}.$ (5.12)
We give just the final result for $I_{-}$ in Eq. (5.12),
$I_{-}\approx\frac{\mu B(E-\mathfrak{E})e^{\mathrm{i}(E-\mathfrak{E})L}}{4\pi\mathfrak{E}L}\left(1-\frac{(\mu B)^{2}}{2E\mathfrak{E}}\right).$ (5.13)
As we mentioned in Sec. 4, the contributions of the diagonal propagators $\Sigma_{aa}$ to the matrix element $\mathcal{M}_{e\to\bar{\mu}}$ are vanishing.
Since $\Sigma_{21}=-\Sigma_{12}$, one gets from Eqs. (3.11), (5.10), and (5.13) that
$\mathcal{M}_{e\to\bar{\mu}}=I_{-}-I_{+}=-\frac{\mathrm{i}Ee^{\mathrm{i}EL}}{2\pi L}\frac{\mu B}{\mathfrak{E}}\left(1-\frac{(\mu B)^{2}}{2\mathfrak{E}E}\right)\left(\sin\mathfrak{E}L-\mathrm{i}\frac{\mathfrak{E}}{E}\cos\mathfrak{E}L\right).$ (5.14)
The probability of the process $\nu_{e}\to\bar{\nu}_{\mu}$ is $P_{\nu_{e}\to\bar{\nu}_{\mu}}\propto|\mathcal{M}_{e\to\bar{\mu}}|^{2}$. Based on Eq. (5.14), we obtain that
$P_{\nu_{e}\to\bar{\nu}_{\mu}}(L)\approx P_{\mathrm{max}}\sin^{2}\left(\sqrt{(\mu B)^{2}+\left(\frac{\Delta m^{2}}{4E}\right)^{2}}L\right),$ (5.15)
where the amplitude of the transition probability has two terms, $P_{\mathrm{max}}=P_{\mathrm{max}}^{(\mathrm{qm})}+P_{\mathrm{max}}^{(\mathrm{qft})}$. These terms have the following forms:
$P_{\mathrm{max}}^{(\mathrm{qm})}=\frac{(\mu B)^{2}}{(\mu B)^{2}+\left(\frac{\Delta m^{2}}{4E}\right)^{2}},\quad P_{\mathrm{max}}^{(\mathrm{qft})}=-\frac{(\mu B)^{4}}{E\left[(\mu B)^{2}+\left(\frac{\Delta m^{2}}{4E}\right)^{2}\right]^{3/2}}.$ (5.16)
We shall see shortly in Sec. 6 that $P_{\mathrm{max}}^{(\mathrm{qm})}$ in Eq. (5.16) reproduces the result of the quantum mechanical approach. The term $P_{\mathrm{max}}^{(\mathrm{qft})}$ in Eq. (5.16) is the QFT contribution to the transition probability. It arises from the terms (i) and (ii) in the propagator $\Sigma_{12}$
To derive Eqs. (5.15) and (5.16), we take into account Eq. (5.7) and assume that $\mu B\sim\mathfrak{E}$. Moreover, we neglect the terms $\sim(\mathfrak{E}/E)^{2}\ll 1$, whereas we keep the terms $\sim\mu B/E$ or $\sim\mathfrak{E}/E$.
One can see in Eq. (5.16) that the QFT contribution to the transition probability is negative. Moreover, $|P_{\mathrm{max}}^{(\mathrm{qft})}|/P_{\mathrm{max}}^{(\mathrm{qm})}\sim\mu B/E\sim\mathfrak{E}/E$. Thus, the QFT term is much smaller than the quantum mechanical one for all realistic oscillations channels where neutrinos are ultrarelativistic.
The latter result proves the assumption that the terms (i) and (ii) in $\Sigma_{12}$ in Eq. (4.5) can be safely neglected for ultrarelativistic neutrinos since their contribution to the transition probability is small. We recall that all the terms in $\Sigma_{12}$ in Eq. (4.5), including (i) and (ii), are of the same order of magnitude. Therefore, before the transition probability computation, one cannot say whether certain terms in $\Sigma_{12}$ are smaller or greater.
5.1 Scattering of charged leptons with arbitrary energies
We have just demonstrated that the contributions of the terms (i) and (ii) in the propagator in Eq. (4.5), which have the QFT origin, to the transition probability of the spin-flavor precession are small. The washing out of these terms happens for ultrarelativistic neutrinos. That is why, we can neglect the terms (i) and (ii) while we study the case of charged leptons with arbitrary energies, with neutrinos being ultrarelativistic.
For simplicity, we take that an incoming electron still propagates along $\mathbf{L}$. The outgoing antimuon can have an arbitrary direction of its momentum. The spin of an electron is taken to be opposite to the electron momentum, $\mathbf{s}_{e}\downuparrows\mathbf{p}_{e}$, whereas the antimuon spin is supposed to be along the particle momentum, $\mathbf{s}_{\bar{\mu}}\upuparrows\mathbf{p}_{\bar{\mu}}$.
The chiral projections of leptons bispinors have the forms,
$\kappa_{e\mathrm{L}}=\sqrt{\frac{1+v_{e}}{2}}\left(\begin{array}[]{c}0\\ 1\end{array}\right),\quad\kappa_{\bar{\mu}\mathrm{R}}=-\sqrt{\frac{1+v_{\bar{\mu}}}{2}}\left(\begin{array}[]{c}e^{-\mathrm{i}\phi_{\bar{\mu}}}\cos\frac{\theta_{\bar{\mu}}}{2}\\ e^{\mathrm{i}\phi_{\bar{\mu}}}\sin\frac{\theta_{\bar{\mu}}}{2}\end{array}\right),$ (5.17)
where $v_{e}$ and $v_{\bar{\mu}}$ are the velocities of the electron and the antimuon, and the spherical angles $\phi_{\bar{\mu}}$ and $\theta_{\bar{\mu}}$ fix the direction of the antimuon momentum with respect to $\mathbf{e}_{z}$.
The averaging of $\Sigma_{12}$ over $\kappa_{\bar{\mu}\mathrm{R}}$ and $\kappa_{e\mathrm{L}}$ is similar to that in Eq. (5.1) except the following replacements:
$f(\rho,\phi,z)\to$ $\sqrt{(1+v_{e})(1+v_{\bar{\mu}})}e^{\mathrm{i}\phi_{\bar{\mu}}}\cos\left(\frac{\theta_{\bar{\mu}}}{2}\right)\frac{\mu B}{4}\left(1+\frac{z}{\sqrt{\rho^{2}+z^{2}}}\right),$
$\mathcal{E}_{\pm}(\mathbf{q})\to$ $\bar{E}\pm\sqrt{\Delta E^{2}+(\mu B)^{2}}.$ (5.18)
The subsequent computation of the matrix element is identical to that in Sec. 5. The final result reads
$\mathcal{M}_{e\to\bar{\mu}}=\frac{\mathrm{i}Ee^{\mathrm{i}EL}}{4\pi L}\sqrt{(1+v_{e})(1+v_{\bar{\mu}})}e^{\mathrm{i}\phi_{\bar{\mu}}}\cos\left(\frac{\theta_{\bar{\mu}}}{2}\right)\frac{\mu B}{\mathfrak{E}}\sin\mathfrak{E}L,$ (5.19)
i.e. we remove the QFT corrections from Eq. (5.14) and take into account the additional factor $e^{\mathrm{i}\phi_{\bar{\mu}}}\cos\left(\tfrac{\theta_{\bar{\mu}}}{2}\right)\sqrt{(1+v_{e})(1+v_{\bar{\mu}})}/2$.
The measurable quantity of the process $e+N\to\bar{\mu}+\tilde{N}^{\prime}$ is the differential cross-section [25, p. 252],
$\mathrm{d}\sigma_{e\to\bar{\mu}}=G_{\mathrm{int}}^{4}E^{2}|\mathcal{M}_{e\to\bar{\mu}}|^{2}\frac{v_{\bar{\mu}}}{v_{e}}\mathrm{d}\Omega_{\bar{\mu}},$ (5.20)
where $\mathrm{d}\Omega_{\bar{\mu}}=2\pi\sin\theta_{\bar{\mu}}\mathrm{d}\theta_{\bar{\mu}}$ is the solid angle of the antimuon momentum space.
Integrating Eq. (5.20) over $\mathrm{d}\Omega_{\bar{\mu}}$, one gets that the total cross-section is
$\sigma_{e\to\bar{\mu}}=\int\mathrm{d}\sigma_{e\to\bar{\mu}}=\frac{E^{4}G_{\mathrm{int}}^{4}}{2\pi L^{2}}P^{(\mathrm{eff})}_{\nu_{e}\to\bar{\nu}_{\mu}},\quad P^{(\mathrm{eff})}_{\nu_{e}\to\bar{\nu}_{\mu}}=(1+v_{e})(1+v_{\bar{\mu}})\frac{v_{\bar{\mu}}}{4v_{e}}\left(\frac{\mu B}{\mathfrak{E}}\right)^{2}\sin^{2}(\mathfrak{E}L).$ (5.21)
One can see in Eq. (5.21) that the total cross-section oscillates in space and is proportional to the quantum mechanical transition probability, $P_{\mathrm{max}}^{(\mathrm{qm})}\sin^{2}(\mathfrak{E}L)$; cf. Eqs. (5.15) and (5.16).
For ultrarelativistic charged leptons, $P^{(\mathrm{eff})}_{\nu_{e}\to\bar{\nu}_{\mu}}\approx P^{(\mathrm{qm})}_{\nu_{e}\to\bar{\nu}_{\mu}}$ since $v_{\bar{\mu}}\sim v_{e}\sim 1$. In general situation, $P^{(\mathrm{eff})}_{\nu_{e}\to\bar{\nu}_{\mu}}<P^{(\mathrm{qm})}_{\nu_{e}\to\bar{\nu}_{\mu}}$ since $\mathfrak{m}_{\mu}>\mathfrak{m}_{e}$. Thus, the factor $(1+v_{e})(1+v_{\bar{\mu}})v_{\bar{\mu}}/(4v_{e})<1$ should be accounted for while designing a neutrino oscillations experiment.
6 Quantum mechanical description of the spin-flavor precession
In this section, we analyze the quantum mechanical description of the spin-flavor precession of Majorana neutrinos in an external magnetic field. This study can be done in both flavor and mass bases. We take the flavor basis since it is more common in the standard quantum mechanical approach.
We suppose that two flavor neutrinos, $\nu_{e}$ and $\nu_{\mu}$, propagate along the $z$-axis and interact with the transverse external magnetic field $\mathbf{B}=(B,0,0)$, as in Sec. 4. Despite one can reveal the nature of neutrinos only in the mass basis, we can still assume that $\nu_{\lambda}$, $\lambda=e,\mu$, are Majorana particles. The interaction of these neutrinos with the magnetic field is owing to the transition magnetic moment $\mu$. The effective Schrödinger equation in the basis $\nu^{\mathrm{T}}=(\nu_{e},\nu_{\mu},\bar{\nu}_{e},\bar{\nu}_{\mu})$ was obtained in Ref. [7],
$\mathrm{i}\dot{\nu}=H_{f}\nu,\quad H_{f}=\left(\begin{array}[]{cccc}-\frac{\Delta m^{2}}{4E_{\nu}}\cos 2\theta&\frac{\Delta m^{2}}{4E_{\nu}}\sin 2\theta&0&\mathrm{i}\mu B\\ \frac{\Delta m^{2}}{4E_{\nu}}\sin 2\theta&\frac{\Delta m^{2}}{4E_{\nu}}\cos 2\theta&-\mathrm{i}\mu B&0\\ 0&\mathrm{i}\mu B&-\frac{\Delta m^{2}}{4E_{\nu}}\cos 2\theta&\frac{\Delta m^{2}}{4E_{\nu}}\sin 2\theta\\ -\mathrm{i}\mu B&0&\frac{\Delta m^{2}}{4E_{\nu}}\sin 2\theta&\frac{\Delta m^{2}}{4E_{\nu}}\cos 2\theta\end{array}\right),$ (6.1)
where $E_{\nu}$ is the mean energy of neutrinos, and $\theta$ is the vacuum mixing angle defined in Eq. (2.1). We already mentioned in Sec. 2 (see also Ref. [22]) that the magnetic moments matrix $(\mu_{ab})=\mu\sigma_{2}$ is hermitian and purely imaginary in the mass basis. Taking into account Eq. (2.1), we get that this matrix in the flavor basis reads $(M_{\lambda\lambda^{\prime}})=U(\mu_{ab})U^{\dagger}=\mu\sigma_{2}$, that is used in Eq. (6.1).
The general solution of Eq. (6.1), which satisfies the given initial condition $\nu_{0}$, has the form,
$\nu(t)=\left(\left[\left(U_{1}\otimes U_{1}^{\dagger}\right)+\left(U_{2}\otimes U_{2}^{\dagger}\right)\right]e^{-\mathrm{i}\mathfrak{\mathfrak{E}_{\nu}}t}+\left[\left(V_{1}\otimes V_{1}^{\dagger}\right)+\left(V_{2}\otimes V_{2}^{\dagger}\right)\right]e^{\mathrm{i}\mathfrak{E}_{\nu}t}\right)\nu_{0},$ (6.2)
where
$\mathfrak{E}_{\nu}=\sqrt{\left(\frac{\Delta m^{2}}{4E_{\nu}}\right)^{2}+(\mu B)^{2}},$ (6.3)
and
$U_{1}$ $=\frac{1}{\sqrt{2\mathfrak{E}_{\nu}}}\left(\begin{array}[]{c}\frac{\mathrm{i}\mu B}{\sqrt{\mathfrak{E}_{\nu}+\frac{\Delta m^{2}}{4E_{\nu}}\cos 2\theta}}\\ 0\\ \frac{\frac{\Delta m^{2}}{4E_{\nu}}\sin 2\theta}{\sqrt{\mathfrak{E}_{\nu}+\frac{\Delta m^{2}}{4E_{\nu}}\cos 2\theta}}\\ \sqrt{\mathfrak{E}_{\nu}+\frac{\Delta m^{2}}{4E_{\nu}}\cos 2\theta}\end{array}\right),\quad V_{1}=\frac{1}{\sqrt{2\mathfrak{E}_{\nu}}}\left(\begin{array}[]{c}\sqrt{\mathfrak{E}_{\nu}+\frac{\Delta m^{2}}{4E_{\nu}}\cos 2\theta}\\ -\frac{\frac{\Delta m^{2}}{4E_{\nu}}\sin 2\theta}{\sqrt{\mathfrak{E}_{\nu}+\frac{\Delta m^{2}}{4E_{\nu}}\cos 2\theta}}\\ 0\\ \frac{\mathrm{i}\mu B}{\sqrt{\mathfrak{E}_{\nu}+\frac{\Delta m^{2}}{4E_{\nu}}\cos 2\theta}}\end{array}\right),$ (6.12)
$U_{2}$ $=\frac{1}{\sqrt{2\mathfrak{E}_{\nu}}}\left(\begin{array}[]{c}\frac{\frac{\Delta m^{2}}{4E_{\nu}}\sin 2\theta}{\sqrt{\mathfrak{E}_{\nu}+\frac{\Delta m^{2}}{4E_{\nu}}\cos 2\theta}}\\ \sqrt{\mathfrak{E}_{\nu}+\frac{\Delta m^{2}}{4E_{\nu}}\cos 2\theta}\\ \frac{\mathrm{i}\mu B}{\sqrt{\mathfrak{E}_{\nu}+\frac{\Delta m^{2}}{4E_{\nu}}\cos 2\theta}}\\ 0\end{array}\right),\quad V_{2}=\frac{1}{\sqrt{2\mathfrak{E}_{\nu}}}\left(\begin{array}[]{c}0\\ \frac{\mathrm{i}\mu B}{\sqrt{\mathfrak{E}_{\nu}+\frac{\Delta m^{2}}{4E_{\nu}}\cos 2\theta}}\\ \sqrt{\mathfrak{E}_{\nu}+\frac{\Delta m^{2}}{4E_{\nu}}\cos 2\theta}\\ -\frac{\frac{\Delta m^{2}}{4E_{\nu}}\sin 2\theta}{\sqrt{\mathfrak{E}_{\nu}+\frac{\Delta m^{2}}{4E_{\nu}}\cos 2\theta}}\end{array}\right),$ (6.21)
are the eigenvectors which are normalized, $|U_{1,2}|^{2}=|V_{1,2}|^{2}=1$.
We suppose that $\nu(0)\equiv\nu_{e}$ initially. Thus, we take $\nu_{0}^{\mathrm{T}}=(1,0,0,0)$ in Eq. (6.2). Using Eq. (6.12), the general solution in Eq. (6.2), satisfying the given initial condition, takes the form,
$\nu(t)=\frac{1}{2\mathfrak{E}_{\nu}}\left(\begin{array}[]{c}\frac{(\mu B)^{2}+\left(\frac{\Delta m^{2}}{4E_{\nu}}\right)^{2}\sin^{2}2\theta}{\mathfrak{E}_{\nu}+\frac{\Delta m^{2}}{4E_{\nu}}\cos 2\theta}e^{-\mathrm{i}\mathfrak{E}_{\nu}t}+\left(\mathfrak{E}_{\nu}+\frac{\Delta m^{2}}{4E_{\nu}}\cos 2\theta\right)e^{\mathrm{i}\mathfrak{E}_{\nu}t}\\ -\mathrm{i}\frac{\Delta m^{2}}{2E_{\nu}}\sin 2\theta\sin\mathfrak{E}_{\nu}t\\ 0\\ -2\mu B\sin\mathfrak{E}_{\nu}t\end{array}\right),$ (6.22)
Based on Eq. (6.22), we get that $\bar{\nu}_{\mu}(t)=-\frac{\mu B}{\mathfrak{E}_{\nu}}\sin\mathfrak{E}_{\nu}t$. The transition probability for the process $\nu_{e}\to\bar{\nu}_{\mu}$ is
$P_{\nu_{e}\to\bar{\nu}_{\mu}}(t)=|\bar{\nu}_{\mu}(t)|^{2}=\frac{(\mu B)^{2}}{(\mu B)^{2}+\left(\frac{\Delta m^{2}}{4E_{\nu}}\right)^{2}}\sin^{2}\left(\sqrt{(\mu B)^{2}+\left(\frac{\Delta m^{2}}{4E_{\nu}}\right)^{2}}t\right).$ (6.23)
It is interesting to mention that the transition probability in Eq. (6.23) does not depend on the vacuum mixing angle.
Now, we can compare the quantum mechanical and quantum field theory approaches to the neutrino spin-flavor precession. If neutrinos are ultrareletivistic, we can put $E_{\nu}=E$, $\mathfrak{E}_{\nu}=\mathfrak{E}$, and $t\approx L$ in Eq. (6.23). Then, we omit the small correction $P_{\mathrm{max}}^{(\mathrm{qft})}$ in Eq. (5.16). Under these approximations, the transition probability in Eqs. (5.15) and (5.16) coincides with that in Eq. (6.23), i.e. both approaches provide the same description of the spin-flavor precession.
7 Discussion
In the present work, we have developed the QFT description of neutrino the spin-flavor precession in a magnetic field. We were based on the QFT formalism, which was originally developed in Refs. [14, 15, 16], where neutrinos are virtual particles. This method was used in Ref. [14] to study neutrino-to-antineutrino oscillations in vacuum. If one deals with Majorana neutrinos with transition magnetic moments, $\nu\leftrightarrow\bar{\nu}$ oscillations are analogous to spin-flavor transitions in a magnetic field. We have adapted the results of Ref. [14] for the description of the spin-flavor precession.
In Sec. 2, we have reminded how Majorana neutrinos can interact with electromagnetic fields. Considering the system of two neutrinos, we have obtained the wave equations for neutrino mass eigenstates in an external magnetic field. The magnetic interaction was shown to mix different mass eigenstates, as well as induce particle-to-antiparticle transitions.
The basics of the QFT approach for the neutrino spin-flavor precession have been discussed in Sec. 3. We have obtained the matrix element corresponding to the transition $\nu_{\beta}\to\bar{\nu}_{\alpha}$ in an external magnetic field. The dressed propagators for the neutrino mass eigenstates, accounting for the magnetic interaction, have been derived in Sec. 4 for neutrinos with arbitrary energies. Then, we have considered the case of ultrarelativistic particles.
Finally, we have derived the transition probability for the $\nu_{e}\to\bar{\nu}_{\mu}$ precession in a magnetic field in Sec. 5 within the QFT approach. This transition probability has the leading term, which was shown in Sec. 6 to coincide with the result of the standard quantum mechanical approach. Moreover, we have obtained that the dressed propagator in Eq. (4.5) acquires the terms (i) and (ii) which have the QFT nature. These terms cannot be neglected before the transition probability calculation. We have demonstrated in Sec. 5 that the QFT terms in the propagator result in the correction to the leading term in the transition probability which is small if neutrinos are ultrarelativistic; cf. Eqs. (5.15) and (5.16).
In Sec. 5.1, we have discussed a realistic situation when charged leptons have arbitrary energies. The forward scattering approximation is not valid in this case. We have derived the effective transition probability in Eq. (5.21) which depends on the velocities of incoming and outgoing charged leptons. It turns out that $P^{(\mathrm{eff})}_{\nu_{e}\to\bar{\nu}_{\mu}}<P^{(\mathrm{qm})}_{\nu_{e}\to\bar{\nu}_{\mu}}$.
One of the main goals of the present work is methodological. It is aimed to demonstrate that the neutrino spin-flavor precession can be described purely with help of the standard QFT methods if reasonable approximation are made.
The major assumption is the ultrarelativity of particles involved in the system. First, in Sec. 5, we assume that charged leptons are utrarelativistic, i.e. $E_{\alpha,\beta}\gg\mathfrak{m}_{\alpha,\beta}$. The assumption of ultrarelativistic charged leptons was used to derive the transition probability for the spin-flavor precession in the forward scattering approximation. We considered a general situation of charged leptons with arbitrary energies in Sec. 5.1. Thus, the assumption of ultrarelativistic charged leptons is not crucial for the results of the work.
The most important approximation is the ultrarelativity of virtual neutrinos. It can be expressed in the form, $E=(E_{\alpha}+E_{\beta})/2\gg m_{1,2}$. Along with the spin-flavor precession under the influence of a magnetic field, for Majorana particles, one has neutrino-to-antineutrino transitions owing to a nonzero neutrino mass. Thus, the series in Eq. (4), would have a more complicated form. Moreover, the bare propagators, derived in Appendix, have rather simple form for ultrarelativistic neutrinos. Finally, without the assumption of the neutrinos ultrarelativity, the computation of the contour integrals, $I_{+}$ and $I_{-}$, would be impossible in Sec. 5. For neutrinos with arbitrary energies, the equation $E-\mathcal{E}_{\pm}(\mathbf{q})+\mathrm{i}0=0$, which defines the poles, has eight roots which should found analytically. Thus, the assumption of ultrarelativistic neutrinos in crucial to treat the spin-flavor precession in frames of QFT.
The next important approximation, which was also mentioned in Ref. [20], is the consideration of two mass eigenstates only. This feature allows one to avoid the branching of the Feynman graphs in Fig. 2.
In deriving of Eq. (5), we made an assumption of the great propagation distance $L$. It allows one to avoid terms $\propto L^{-2}$ in Eq. (5). It is, however, known that the coherence in neutrino oscillations is lost if $L>L_{\mathrm{coh}}$, where $L_{\mathrm{coh}}$ is the coherence length. It is found that $L_{\mathrm{coh}}\propto L_{\mathrm{osc}}\sim\mathfrak{E}^{-1}$ (see, e.g., Ref. [12]).
In fact, numerous factors can contribute to the coherence length. For instance, it depends on a particular experiment where neutrino oscillations are studied (see, e.g., Ref. [26]). A careful determination of $L_{\mathrm{coh}}$ requires the replacement of the plane waves wavefunctions of charged leptons with wave packets of a finite width $\sigma_{x}$ [15, 26]. This approach is applicable for the studies of neutrino oscillations in vacuum in frames of QFT. In that situation, the expression for the matrix element can be derived analytically for plane waves wavefunctions of charged leptons, with both neutrinos and charged leptons having arbitrary energies.
In our case, when we study neutrinos interacting with an external magnetic field, the integrations in Sec. 5 are quite nontrivial even for plane waves charged leptons. As we mentioned earlier, the matrix element can be derived only in the approximation of ultrarelativistic neutrinos. Therefore, a careful consideration of the decoherence in the spin-flavor precession in frames of QFT is a quite complicated task which is beyond the scope of the present work. We can just roughly estimate the coherence length as $L_{\mathrm{coh}}\sim\sigma_{x}E/\mathfrak{E}$ for ultrarelativistic neutrinos [15]. Another estimate for the coherence length in a magnetic field reads [27] $L_{\mathrm{coh}}\sim\sigma_{x}E^{3}/\mathfrak{E}m_{a}^{2}$.
We also make an interesting observation about neutrino oscillations in external fields. The neutrino interaction with background matter or with a magnetic field can be rather trivially incorporated to the neutrino vacuum oscillations description. Indeed, one just adds a linear term with a certain external field to the vacuum effective Hamiltonian for flavor neutrinos. However, from the point of view of QFT, the effect of external fields on neutrino oscillations is an essentially nonperturbative phenomenon. The correct form of the transition probability is obtained only if all terms in the infinite series in Eq. (4) are summed up.
Acknowledgments
I am thankful to M. Deka for the help with the Feynman diagram drawing.
Appendix A Propagators of Weyl neutrinos in vacuum
In this Appendix, we study the propagators of Majorana neutrinos in vacuum. Using the Weyl form of the neutrino wavefunctions, we represent the propagator for left-handed neutrinos and derive the propagator for right-handed antineutrinos.
The propagator of the field $\eta_{a}$, which corresponds to left-handed neutrinos, was obtained in Ref. [28] (see also Ref. [20], where this propagator was derived in the presence of background matter). Using this result, we represent the propagator in question in the form,
$S_{a}^{(\eta)}(x)=\mathrm{i}\int\frac{d^{4}p}{(2\pi)^{4}}e^{-\mathrm{i}px}\frac{1}{2}\left\{\left[1-\frac{(\bm{\sigma}\mathbf{p})}{E_{a}}\right]\frac{1}{p_{0}-E_{a}+\mathrm{i}0}+\left[1+\frac{(\bm{\sigma}\mathbf{p})}{E_{a}}\right]\frac{1}{p_{0}+E_{a}-\mathrm{i}0}\right\},$ (A.1)
where $E_{a}=\sqrt{\mathbf{p}^{2}+m_{a}^{2}}$ and $\mathrm{i}0$ stays for a small imaginary term. We can decompose Eq. (A.1) as $S_{a}^{(\eta)}=S_{a\mathrm{L}}+\tilde{S}_{a\mathrm{R}}$, where
$S_{a\mathrm{L}}(p)=\frac{\mathrm{i}}{2}\left[1-\frac{(\bm{\sigma}\mathbf{p})}{E_{a}}\right]\frac{1}{p_{0}-E_{a}+\mathrm{i}0},$ (A.2)
is the propagator of left-handed neutrinos which is used in Sec. 4.
The the wavefunction $\xi_{a}$, corresponding to a right handed antineutrino, obeys the equation,
$\mathrm{i}\dot{\xi}_{a}-(\bm{\sigma}\mathbf{p})\xi_{a}-\mathrm{i}m_{a}\sigma_{2}\xi_{a}^{*}=0,$ (A.3)
which can be obtained from the Dirac equation for a Majorana neutrino if we take that the bispinor has the form, $\psi_{a}^{\mathrm{T}}=(\xi_{a},-\mathrm{i}\sigma_{2}\xi_{a}^{*})$. Such a bispinor satisfies the Majorana condition, $(\psi_{a})^{c}=\psi_{a}$. The general solution of Eq. (A.3) has the form,
$\xi_{a}(\mathbf{x},t)=$ $\int\frac{\mathrm{d}^{3}p}{(2\pi)^{3/2}}\sqrt{\frac{E_{a}+p}{E_{a}}}\Big{[}\left(a_{a+}w_{+}+A_{a-}a_{a-}w_{-}\right)e^{-\mathrm{i}E_{a}t+\mathrm{i}\mathbf{px}}$
$+\left(a_{a-}^{\dagger}w_{+}+A_{a+}a_{a+}^{\dagger}w_{-}\right)e^{\mathrm{i}E_{a}t-\mathrm{i}\mathbf{px}}\Big{]},$ (A.4)
where
$A_{a\pm}=\pm\frac{m_{a}}{E_{a}+p},$ (A.5)
and
$w_{+}=\left(\begin{array}[]{c}e^{-\mathrm{i}\varphi/2}\cos\vartheta/2\\ e^{\mathrm{i}\varphi/2}\sin\vartheta/2\end{array}\right),\quad w_{-}=\left(\begin{array}[]{c}-e^{-\mathrm{i}\varphi/2}\sin\vartheta/2\\ e^{\mathrm{i}\varphi/2}\cos\vartheta/2\end{array}\right),$ (A.6)
are the helicity amplitudes, which satisfy the relation, $w_{\pm}\otimes w_{\pm}^{\dagger}=\frac{1}{2}\left[1\pm\bm{\sigma}\hat{p}\right$]. In Eq. (A.6), the angles $\vartheta$ and $\varphi$ fix the direction of the neutrino momentum $\mathbf{p}$. The creation and annihilation operators, $a_{a\pm}^{\dagger}(\mathbf{p})$ and $a_{a\pm}(\mathbf{p})$, obey the anticommurators,
$\left\{a_{a\pm}(\mathbf{p}),a_{b\pm}^{\dagger}(\mathbf{q})\right\}_{+}=\delta_{ab}\delta(\mathbf{p}-\mathbf{q}),$ (A.7)
with the rest of the anticommutators being equal to zero.
Then, we define the propagator of $\xi_{a}$ as
$S_{a}^{(\xi)}(x-y)=\theta(x_{0}-y_{0})\left\langle 0\right|\xi(x)\xi^{\dagger}(y)\left|0\right\rangle-\theta(y_{0}-x_{0})\left\langle 0\right|\xi^{*}(y)\xi^{\mathrm{T}}(x)\left|0\right\rangle.$ (A.8)
Using Eqs. (A)-(A.7) and making calculation analogous to those in Refs. [28, 20], we cast Eq. (A.8) to the form,
$S_{a}^{(\xi)}(x)=\mathrm{i}\int\frac{d^{4}p}{(2\pi)^{4}}e^{-\mathrm{i}px}\frac{1}{2}\left\{\left[1+\frac{(\bm{\sigma}\mathbf{p})}{E_{a}}\right]\frac{1}{p_{0}-E_{a}+\mathrm{i}0}+\left[1-\frac{(\bm{\sigma}\mathbf{p})}{E_{a}}\right]\frac{1}{p_{0}+E_{a}-\mathrm{i}0}\right\}.$ (A.9)
Similarly to $S_{a}^{(\eta)}$ in Eq. (A.1), we decompose $S_{a}^{(\xi)}$ in Eq. (A.9) as $S_{a}^{(\xi)}=S_{a\mathrm{R}}+\tilde{S}_{a\mathrm{L}}$, where
$S_{a\mathrm{R}}(p)=\frac{\mathrm{i}}{2}\left[1+\frac{(\bm{\sigma}\mathbf{p})}{E_{a}}\right]\frac{1}{p_{0}-E_{a}+\mathrm{i}0},$ (A.10)
is the propagator of right-handed antineutrinos which is used in Sec. 4.